### v1

# Provably Finding a Hidden Dense Submatrix among Many Planted Dense Submatrices via Convex Programming

Abstract
We consider the densest submatrix problem, which seeks the submatrix of fixed size of a given binary matrix that contains the most nonzero entries. This problem is a natural generalization of fundamental problems in combinatorial optimization, e.g., the densest subgraph, maximum clique, and maximum edge biclique problems, and has wide application the study of complex networks. Much recent research has focused on the development of sufficient conditions for exact solution of the densest submatrix problem via convex relaxation. The vast majority of these sufficient conditions establish identification of the densest submatrix within a graph containing exactly one large dense submatrix hidden by noise. The assumptions of these underlying models are not observed in real-world networks, where the data may correspond to a matrix containing many dense submatrices of varying sizes.
We extend and generalize these results to the more realistic setting where the input matrix may contain many large dense subgraphs. Specifically, we establish sufficient conditions under which we can expect to solve the densest submatrix problem in polynomial time for random input matrices sampled from a generalization of the stochastic block model. Moreover, we also provide sufficient conditions for perfect recovery under a deterministic adversarial. Numerical experiments involving randomly generated problem instances and real-world collaboration and communication networks are used empirically to verify the theoretical phase-transitions to perfect recovery given by these sufficient conditions.
1 Introduction
We consider the densest matrix problem (DSM), in which one aims to identify the submatrix of given size containing a maximum number of nonzero entries; see Lanciano et al. (2024) for a survey of the historical and recent study of the DSM. The densest submatrix can be considered a generalization of the densest bipartite subgraph problem: given a bipartite graph $G=(V,U,E)$ and integers $k_{1}$ and $k_{2}$, the objective is to extract a subgraph induced by $k_{1}$ nodes from $V$ and $k_{2}$ nodes from $U$ that maximizes the number of edges. This further generalizes the densest subgraph problem: given graph $G$ and integer $k$, find the $k$-node subgraph containing maximum number of edges.
The problem of discovering dense subgraphs has garnered significant attention due to its broad applicability in domains such as bioinformatics Malod-Dognin et al. (2010); Gómez Ravetti and Moscato (2008), financial and economic network analysis Boginski et al. (2006), social network mining Balasundaram et al. (2011); Pattillo et al. (2011), telecommunications Jain et al. (2003), clustering Schaeffer (2007), computational biology, and image processing. In many such contexts, densely connected substructures often correspond to functionally meaningful modules, such as communities in social graphs or coherent biological or transactional clusters.
From a computational standpoint, the densest subgraph problem of fixed size is known to be NP-hard Karp (2009) and is considered difficult to approximate within any constant factor in general graphs Andersen (2007). This intractability arises from its close relationship to the maximum clique problem, a classical NP-hard problem. In fact, it has been shown that unless certain complexity-theoretic assumptions fail, no polynomial time approximation schemes (PTAS) exist for the densest $k$-subgraph problem Feige (2002); Khot (2006); Alon et al. (1998, 2011).
Despite this worst-case intractability, there is a growing body of literature establishing sufficient conditions for exact solution of the densest submatrix problem, or special cases such as the maximum clique and densest subgraph problem, via spectral methods and/or convex relaxation Ames (2015); Candès et al. (2011); Chen and Xu (2014); Doan and Vavasis (2013); Ames and Vavasis (2011); Feige and Krauthgamer (2000); Bombina and Ames (2020); Sohn and Wein (2025); Dadon et al. (2024); Omer and Ali (2023, 2025). Most of these results establish that if the given problem instance corresponds to a single dense submatrix hidden by sparse noise, then this hidden submatrix is the exact solution of some convex optimization problem.
Our proposed convex relaxation adopts a nuclear norm minimization strategy, analogous to that used in robust PCA Chandrasekaran et al. (2011), to capture low-rank structure while tolerating sparse noise. This adaptation enables recovery guarantees in planted submatrix settings that generalize and improve upon earlier works. For example, Ames and Vavasis Ames and Vavasis (2011) proposed a nuclear norm relaxation for the maximum-edge biclique problem, providing recovery guarantees based on Chernoff bounds and spectral norm analysis. Ames Ames (2015) further improved this approach by combining nuclear norm minimization with an $\ell_{1}$-penalty. Their model decomposes the adjacency matrix into low-rank and sparse components, linking the formulation to Robust Principal Component Analysis (RPCA). Bombina and Ames Bombina and Ames (2020) established tight recovery guarantees for planted dense $m\times n$ submatrices under constant signal assumptions, proving that perfect recovery is possible when $\min\{m,n\}\geq\Omega((\log N)^{3/2})$ under sparse noise. They also demonstrated phase transition phenomena for convex formulations. However, their analysis was restricted to homogeneous or block-constant models and did not account for general heterogeneous structures where each block has its own distinct probability distribution.
A key common factor of these earlier results is the foundational assumption that the input matrix contains a single dense hidden submatrix. In contrast, we generalize these results to a more realistic setting where the input matrix may contain many dense submatrices. Specifically, we consider a heterogeneous rectangular stochastic block model, where each block is governed by its Bernoulli probability distribution. Specifically, the observed binary matrix $\boldsymbol{A}\in\{0,1\}^{M\times N}$ is constructed by partitioning the row indices $\{1,\dots,M\}$ into $k_{1}$ disjoint subsets $U_{1},\dots,U_{k_{1}}$ and the column indices $\{1,\dots,N\}$ into $k_{2}$ disjoint subsets $V_{1},\dots,V_{k_{2}}$. For each block pair $(U_{r},V_{s})$, the entries $A_{ij}$ with $i\in U_{r}$ and $j\in V_{s}$ are independently sampled from a Bernoulli distribution with parameter $p_{rs}$, defining the local edge probability of that rectangular block.
1.1 Contributions
The primary contribution of our work is a sufficient condition for identification of a planted dense submatrix hidden among potentially many other dense submatrices. We will see that this result can be specialized to provide sufficient conditions for identification of planted cliques, bicliques, and dense subgraphs in graphs containing many such combinatorial objects. In all cases, we will recover the hidden submatrix or subgraph from the solution of a convex relaxation of the densest $m\times n$-submatrix problem, which we will formally define next.
Definition 1.1 (Densest Submatrix Problem).
Given binary matrix $\boldsymbol{A}\in\{0,1\}^{M\times N}$ and integers $m\in\{1,2,\dots,M\}$ and $n\in\{1,2,\dots,N\}$, the densest $m\times n$-submatrix problem seeks the submatrix of $\boldsymbol{A}$ with $m$ rows and $n$ columns that contains the most non-zero entries.
Our main theorem will provide sufficient conditions under which exact recovery of the planted structure is guaranteed with high probability as defined below.
Remark 1 (Interpretation of “with high probability”).
In this work, the phrase “with high probability” (abbreviated as w.h.p.) refers to events whose probability converges to 1 as the problem size increases. Formally, for a sequence of random events $\{E_{n}\}$, we say that $E_{n}$ occurs with high probability if
$\Pr(E_{n})\to 1\quad\text{as }n\to\infty.$
polynomially as $n\to\infty$. Equivalently, $\Pr(E_{n})\geq 1-1/p(n)$ for some polynomial function $p(n)$ such that for any $\varepsilon>0$, there exists an integer $N_{\varepsilon}$ such that for all $n\geq N_{\varepsilon}$, we have $p(n)<\epsilon 0$.
The planted dense matrix problem is NP-Hard by reduction from the maximum clique problem. As such, we should not expect to be able to solve it efficiently in all problem instances. We next define a probabilistic model for random matrices with different expected densities across different blocks. We will use this block structure and these densities to characterize “nice” matrices, for which we should expect to be able to identify their dense submatrices.
Definition 1.2 (Planted Submatrix Model).
Consider partitions $(U_{1},U_{2},\dots,U_{k_{1}})$ and $(V_{1},V_{2},\dots,V_{k_{2}})$ of $\{1,2,\dots,M\}$ and $\{1,2,\dots,N\}$ respectively. We say that matrix $\boldsymbol{A}\in\{0,1\}^{M\times N}$ is sampled from the planted submatrix model corresponding to this partition if each entry of $\boldsymbol{A}$ is independently sampled according to the Bernoulli probability distribution
$\Pr(A_{ij}=1)=p_{rs},\hskip 18.06749pt\Pr(A_{ij}=0)=1-p_{rs}$
for each $i\in U_{r}$, $j\in V_{s}$, $r\in\{1,2,\dots,k_{1}\}$ and $r,s\in\{1,2,\dots,k_{2}\}$.
We next state a simplified version of the planted submatrix model, which we will use to present a simplification of our main recovery guarantee; we will present a more detailed analysis and sufficient conditions for recovery in Section 2.
Definition 1.3 (Balanced Planted Submatrix Model).
We say that square matrix $\boldsymbol{A}\in\{0,1\}^{N\times N}$ is sampled from the balanced planted submatrix model if it is sampled from the planted submatrix model with partitions with square blocks of the same size:
$m_{1}=m_{2}=\cdots=m_{k}=n_{1}=n_{2}=\cdots=n_{k}.$
Note that both the planted submatrix model and the balanced planted submatrix model are generalizations of the stochastic block model (SBM), which is commonly used in the analysis of clustering and community detection algorithms; see Abbe (2023) for a detailed discussion of clustering and community detection under the SBM. Under this model, the binary matrix $\boldsymbol{A}\in\{0,1\}^{M\times N}$ is structured into rectangular blocks, where each block $\boldsymbol{A}(U_{r},V_{s})$ is independently sampled from an Erdös–Rényi distribution with edge probability $p_{rs}$. This results in a heterogeneous block structure in which each block may follow a different connectivity regime.
We will illustrate the planted submatrix and balanced submatrix models in Fig. 1. We sample two matrices, one balanced and the other unbalanced, with $6$ especially dense submatrices in the diagonal blocks. The left panel illustrates the full matrices generated under each model, while the right panel reveals the hidden ground truth structure used in the planting process.
(a) $\boldsymbol{A}$ sampled from the planted submatrix model
(b) $\boldsymbol{B}$ sampled from the balanced planted submatrix model
Figure 1: Examples of matrices sampled from planted submatrix and balanced planted submatrix models with heterogeneous probabilities $\{p_{rs}\}$.
We are now ready to state a simplified version of our key contribution, which provides a sufficient condition on the parameters in the balanced planted submatrix that ensures recovery of the densest submatrix via the solution of a tractable convex program.
Theorem 1.1 (Perfect Recovery in the Balanced Planted Submatrix Model).
Let $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ be sampled from the balanced planted submatrix model with row partitions $(U_{1},\ldots,U_{k})$ and column partitions $(V_{1},\ldots,V_{k})$. Let $m:=|U_{i}|=|V_{i}|$ for all $i=1,2,\dots,k.$ Let $p_{rs}$ denote the probability that $A_{ij}=1$ for $i\in U_{r}$, $j\in V_{s}$.
Assume that $(U_{1},V_{1})$ is the planted block of interest. Define the worst-case variance proxy:
$\tilde{\sigma}_{*}^{2}:=\frac{p^{*}}{1-p^{*}}\text{ where }p^{*}:=\max_{(r,s)\neq(1,1)}\frac{p_{rs}}{1-p_{rs}}.$
Then there exists constant $c>0$ such that
  • the planted block $(U_{1},V_{1})$ is the unique densest $m\times m$ submatrix of $\boldsymbol{A}$ with high probability; and
  • we can identify $(U_{1},V_{1})$ in polynomial-time using the solution of a convex program with high probability
if the following conditions hold
$(\log N)^{3}$ $\leq m^{2}$ (1a)
$p_{11}-p^{*}$ $\geq c\cdot\max\left\{\sqrt{\frac{\tilde{\sigma}_{*}^{2}N\log N}{m^{2}}},\sqrt{\frac{\sigma_{11}^{2}\log N}{m}},\frac{\log N}{m}\right\}.$ (1b)
Condition (1b) serves as a signal-to-noise ratio (SNR) balancing the signal, measured by the gap between $p_{11}$ and $p^{*}$, relative to noise, measured by the variances $\tilde{\sigma}^{2}_{*}$ and $\sigma_{11}^{2}$, normalized by the relative gap in the size of the planted blocks, $m^{2}$, and the size of the matrix given by $N$. If this SNR is sufficiently large, then we can expect to find the densest $m\times m$-submatrix in polynomial-time.
This result generalizes to the unbalanced planted submatrix model, as well as adversarially constructed matrices. We will motivate the relevant convex relaxation and provide these sufficient conditions in following section.
2 The Densest Submatrix Problem
2.1 Formulation and Relaxation
We approach the challenge of the intractability of the general densest $m\times n$-submatrix problem by relaxing the problem as a convex program. While we do not claim that the relaxation yields globally optimal solutions for all input matrices, we demonstrate that for structured matrices generated under the planted submatrix model, our convex formulation is capable of exact recovery. In particular, we show that the proposed model recovers the planted $m\times n$ block even when it is obscured by noise in the form of spurious nonzero entries or missing signal entries elsewhere in the matrix.
We define $M]:=\{1,2,...,M\}$ and $N]:=\{1,2,...,N\}$ for every positive integer $M$ and $N$ respectively. Let $\bar{U}_{1}\subseteq[M]$ and $\bar{V}_{1}\subseteq[N]$ denote the true row and column indices of the planted submatrix. Furthermore, we define vectors $\bar{\boldsymbol{u}}\in\{0,1\}^{M}$ and $\bar{\boldsymbol{v}}\in\{0,1\}^{N}$ as the indicator vectors of the planted row and column sets. The planted submatrix support is given by matrix $\bar{\boldsymbol{X}}=\bar{\boldsymbol{u}}\bar{\boldsymbol{v}}^{\top}\in\{0,1\}^{M\times N}$, e. By construction, $\bar{\boldsymbol{X}}=\bar{\boldsymbol{u}}\bar{\boldsymbol{v}}^{\top}$ is a rank-one binary matrix supported on $\bar{U}_{1}\times\bar{V}_{1}$), such that $\bar{X}_{ij}=1$ if $(i,j)\in\bar{U}_{1}\times\bar{V}_{1}$ and $\bar{X}_{ij}=0$ otherwise.
Suppose that $\Omega$ encodes the set of zero entries in a given matrix $\boldsymbol{A}\in\mathbb{R}^{M\times N}$. Without loss of generality, we may assume that $\boldsymbol{A}$’s entries are binary. If not, we can substitute $\boldsymbol{A}$ with a binary matrix that features the same sparsity pattern, preserving the index set of the densest $m_{1}\times n_{1}$-submatrix. Our objective is to derive a rank-one matrix $\boldsymbol{X}$ with $m_{1}n_{1}$ nonzero entries that minimizes the number of disagreements with $\boldsymbol{A}$ on $\Omega$.
The observed binary matrix $\boldsymbol{A}\in\{0,1\}^{M\times N}$ may differ from $\bar{\boldsymbol{X}}$ due to noise or missing entries. In particular, we assume that the zero entries in $\boldsymbol{A}$ are indexed by a set $\Omega\subseteq[M]\times[N]$, representing locations where the observed matrix disagrees with the ground-truth planted block. We define the disagreement matrix $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ as the projection of $\bar{\boldsymbol{X}}$ onto the set $\Omega$, i.e., a binary matrix supported only on indices where $\bar{\boldsymbol{X}}$ and $\boldsymbol{A}$ disagree due to $A_{ij}=0$ while $\bar{X}_{ij}=1$. This leads to the representation
$\tilde{\boldsymbol{A}}_{\bar{U},\bar{V}}=\bar{\boldsymbol{X}}-P_{\Omega}(\bar{\boldsymbol{X}}),$
where $\tilde{\boldsymbol{A}}_{\bar{U},\bar{V}}$ can be interpreted as a noise-corrected approximation of the true submatrix in $\boldsymbol{A}$. The pair $(\bar{\boldsymbol{X}},\bar{\boldsymbol{Y}})$ constitutes the matrix representation of the recovered submatrix and its disagreement pattern.
We define the density of the submatrix indexed by $(\bar{U}_{1},\bar{V}_{1})$ as:
$d(\bar{\boldsymbol{X}})=\frac{1}{m_{1}n_{1}}\left(m_{1}n_{1}-\|P_{\Omega}(\bar{\boldsymbol{X}})\|_{0}\right),$
where $\|P_{\Omega}(\bar{\boldsymbol{X}})\|_{0}$ counts the number of locations in $\bar{\boldsymbol{X}}$ that lie in $\Omega$, where $\boldsymbol{A}$ reports a zero. Since $P_{\Omega}(\bar{\boldsymbol{X}})$ is binary, its support size can be equivalently written as:
$\|P_{\Omega}(\bar{\boldsymbol{X}})\|_{0}=\sum_{(i,j)\in\Omega}[\bar{\boldsymbol{X}}]_{ij}.$
This allows us to cast the planted submatrix recovery problem as a structured rank-one optimization task with explicit modeling of disagreement against the zero-pattern of the data:
$\begin{array}[]{cl}\displaystyle\min_{\boldsymbol{X},\boldsymbol{Y}\in\{0,1\}^{M\times N}}&\mathrm{Tr}(\boldsymbol{Y}\boldsymbol{1}\boldsymbol{1}^{T})\\ \mathrm{s.t.}&\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=m_{1}n_{1},~P_{\Omega}(\boldsymbol{X}-\boldsymbol{Y})=\boldsymbol{0},~\mathrm{rank}(\boldsymbol{X})=1.\end{array}$ (2)
By penalizing the number of mismatches through the matrix $\boldsymbol{Y}$, the problem (Eq. 2) seeks a rank-one submatrix that minimizes disagreements between $\boldsymbol{A}$ and $\boldsymbol{X}$ on $\Omega$ while preserving the target support size $m_{1}n_{1}$.
Relaxing both binary and rank constraints gives the convex relaxation:
$\begin{array}[]{cl}\min&\lVert\boldsymbol{X}\rVert_{*}+\gamma\mathrm{Tr}(\boldsymbol{Y}\boldsymbol{1}\boldsymbol{1}^{T})\\ \mathrm{s.t.}&\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=m_{1}n_{1},~P_{\Omega}(\boldsymbol{X}-\boldsymbol{Y})=\boldsymbol{0},~\boldsymbol{0}\leq\boldsymbol{X}\leq\boldsymbol{1}\boldsymbol{1}^{T},~\boldsymbol{0}\leq\boldsymbol{Y}.\end{array}$ (3)
In the expression, $\gamma\geq 0$ denotes a regularization parameter, selected to balance between the two objectives. As previously mentioned, it is reasonable to anticipate retrieving the solution of (2) from (3) when matrix $\boldsymbol{A}$ consists of a dense $m_{1}\times n_{1}$ submatrix hidden among perhaps many other dense submatrices under certain assumptions about the size and density of these submatrices. In particular, this program succeeds in recovering the planted block exactly when embedded in a noisy binary matrix under the conditions of Theorem 1.1.
2.2 A Sufficient Condition for Perfect Recovery for Random Matrices
Theorem 1.1 provided a sufficient condition for recovery of planted submatrices in a random graph sampled from the balanced planted submatrix model. We now state the generalization of this sufficient condition for the planted submatrix model with (possibly) unbalanced submatrix sizes.
Theorem 2.1.
Let $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ be sampled from the planted submatrix model with row partitions $(U_{1},\ldots,U_{k_{1}})$ and column partitions $(V_{1},\ldots,V_{k_{2}})$. Let $m_{i}=|U_{i}|$ and $n_{i}=|V_{i}|$, and let $p_{rs}$ denote the probability that $A_{ij}=1$ for $i\in U_{r}$, $j\in V_{s}$.
Assume that $(U_{1},V_{1})$ is the planted block of interest and let $\boldsymbol{u}_{1}$, $\boldsymbol{v}_{1}$ denote the characteristic vectors of $U_{1}$ and $V_{1}$. Define $N^{*}=\max\{M,N\}$, $n_{j}^{*}=\max\{m_{j},n_{j}\}$, $n_{j*}=\min\{m_{j},n_{j}\}$, and the worst-case variance proxy:
$\tilde{\sigma}_{*}^{2}:=\max_{(r,s)\neq(1,1)}\left\{\min\left\{\frac{1-p_{rs}}{p_{rs}},\frac{p_{rs}}{1-p_{rs}}\right\}\right\}.$
Then there exist constants $c_{1},c_{2}>0$ such that the planted block $(U_{1},V_{1})$ is the unique densest $m_{1}\times n_{1}$ submatrix of $\boldsymbol{A}$ and $\boldsymbol{X}^{*}=\boldsymbol{u}_{1}\boldsymbol{v}_{1}^{T}$ unique optimal solution of the convex relaxation (3) if the following conditions hold for $(r,s)\neq(1,1)$
$m_{1}n_{1}$ $\leq m_{r}n_{s},$ (4a)
$(\log N)^{3}$ $\leq\min\{m_{r}^{2},n_{s}^{2}\}$ (4b)
$p_{11}-p_{rs}$ $\geq c_{1}\cdot\max\left\{\sqrt{\frac{\tilde{\sigma}_{*}^{2}N^{*}\log N^{*}}{m_{1}n_{1}}},\sqrt{\frac{\sigma_{11}^{2}\log N^{*}}{n_{1*}}},\frac{\log N^{*}}{n_{1*}}\right\}$ (4c)
and the regularization parameter $\gamma$ in (3) belongs to the interval
$\gamma\in\left(\frac{2}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}},\frac{c_{2}}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}}\right),$
where $p^{*}:=\max_{(r,s)\neq(1,1)}p_{rs}$.
This result shows that recovery depends critically on a separation condition between the planted density $p_{11}$ and the background noise levels $p_{rs}$. The condition balances statistical fluctuations, structural asymmetry, and problem size. When it holds, solving the convex program in (3) results in a solution $\boldsymbol{X}$ whose support exactly matches the planted submatrix. It is easy to see that Theorem 2.1 specializes to Theorem 1.1 when $\boldsymbol{A}$ is sampled from the balanced planted submatrix model.
2.3 A Sufficient Condition for Perfect Recovery for Adversarially Generated Matrices
The sufficient conditions for perfect recovery given in Section 2.2 may be overly conservative. For example, the assumption that the planted submatrix of interest is smaller than all other blocks in the generative model is kept largely to ensure that this is the densest $m_{1}\times n_{1}$-submatrix with high probability if the hypothesis of Theorem 2.1 is satisfied. This simplifies the statement of the sufficient conditions, as well as their proof, but is not necessary, in general, to ensure perfect recovery. In this section, we provide a sufficient condition for recovery that holds for deterministically generated graphs. This condition deterministically encodes the relationship between planted submatrices, which ensures perfect recovery in the random case, i.e., that $(U_{1},V_{1})$ sufficiently more dense than all other $m_{1}\times n_{1}$ subgraphs.
2.3.1 The Adversarial Planted Submatrix Model
We start by describing an adversarial analogue of the planted submatrix model (Definition 1.2):
  • Partition $[M]$ and $[N]$ into $K$ blocks given by $(U_{1},U_{2},\dots,U_{K})$ and $(V_{1},V_{2},\dots,V_{K})$, respectively. Let $m_{i}:=|U_{i}|$ and $n_{i}:=|V_{i}|$ for each $i=1,2,\dots,K$. We’ll associate these with a partition of the rows and columns of binary matrix $\boldsymbol{A}\in\mathbb{R}^{M\times N}$.
  • Pick one block, indexed by $(U_{1},V_{1})$ without loss of generality, and set $A_{ij}=1$ for all $i\in U_{1},j\in V_{1}$.
  • An adversary is then allowed to do the following:
    • Delete, i.e., set equal to 0, up to $(1-\tilde{\delta})m_{1}$ entries in the $j$th column of $\boldsymbol{A}[U_{1},V_{1}]$ for each $j\in V_{1}$. Note that this implies that at least $\tilde{\delta}m_{1}$ entries of each column of $\boldsymbol{A}[U_{1},V_{1}]$ remain equal to $1$.
    • Delete up to $(1-\tilde{\delta})n_{1}$ entries in the $i$th row of $\boldsymbol{A}[U_{1},V_{1}]$ for each $i\in U_{1}$.
    • In total, the adversary deletes $\bar{r}_{11}$ entries in $\boldsymbol{A}[U_{1},V_{1}]$; note that $r_{11}:=m_{1}n_{1}-\bar{r}_{11}$ entries of $\boldsymbol{A}[U_{1},V_{1}]$ are equal to $1$.
    • Add, i.e., set equal to $1$, up to $r_{ii}$ nonzero entries in each diagonal block $\boldsymbol{A}[U_{i},V_{i}]$, $i=2,3,\dots,K$.
    • Add up to $\delta m_{1}$ nonzero entries in the $j$th column of $\boldsymbol{A}[U_{1},V-V_{1}]$ for each $j\in V-V_{1}$, with at most $r_{1}$ nonzero entries in $\boldsymbol{A}[U_{1},V-V_{1}]$.
    • Add up to $\delta n_{1}$ nonzero entries in the $i$th row of $\boldsymbol{A}[U-U_{1},V_{1}]$ for each $i\in U-U_{1}$, with at most $r_{2}$ nonzero entries in $\boldsymbol{A}[U-U_{1},V_{1}]$.
    • Add up to $r_{3}$ nonzero entries among $\{(i,j):i\in U_{q},j\in V_{s},\text{such that }q,s\geq 2,q\neq s\}$.
This construction models a two-person game where an adversary attempts to hide a planted all-ones matrix by a combination of creation of other dense blocks and setting entries within the planted matrix equal to 0.
2.3.2 Perfect Recovery Under the Adversarial Planted Submatrix Model
It is reasonable to assume that we should have perfect recovery of the planted submatrix under the adversarial construction described in Section 2.3.1 if the adversary is allowed to introduce a limited amount of noise in the form of entry additions and deletions. The following theorem provides a characterization of the limits on adversarial noise that ensures that the planted submatrix is recoverable as the unique solution of (3).
Theorem 2.2.
Suppose that matrix $\boldsymbol{A}$ is constructed as described in Section 2.3.1 with planted submatrix indexed by $(U_{1},V_{1})$. Let $\boldsymbol{u}_{1}$ and $\boldsymbol{v}_{1}$ denote the characteristic vectors of $U_{1}$ and $V_{1}$ respectively. Then there exists constant $c$, depending only on $\delta,\tilde{\delta}$ such that if
$2\tilde{\delta}-\delta$ $>1$ (5)
$\max\{r_{1},r_{2},r_{3},r_{22},r_{33},\dots,r_{KK},\bar{r}_{11}\}$ $\leq cm_{1}n_{1}$ (6)
then $(U_{1},V_{1})$ indexes the densest $m_{1}\times n_{1}$-submatrix of $\boldsymbol{A}$ and $(\boldsymbol{u}_{1}\boldsymbol{v}_{1}^{T},P_{\Omega}(\boldsymbol{u}_{1}\boldsymbol{v}_{1}^{T}))$ is the unique optimal solution of (3) with
$\gamma=\frac{1}{2\tilde{\delta}-\delta-1}\frac{1}{\sqrt{m_{1}n_{1}}}.$ (7)
Essentially Theorem 2.2 indicates that we have perfect recovery of $\boldsymbol{A}[U_{1},V_{1}]$, unless the adversary is able to create a denser $m_{1}\times n_{1}$-submatrix through a combination of
  • addition of $\Omega(m_{1}n_{1})$ nonzero entries outside $\boldsymbol{A}[U_{1},V_{1}]$;
  • deletion of $\Omega(m_{1}n_{1})$ nonzero entries within $\boldsymbol{A}[U_{1},V_{1}]$; or
  • adding $\Omega(m_{1})$ entries within $\boldsymbol{A}[U_{1},V-V_{1}]$ and/or $\Omega(n_{1})$ entries within $\boldsymbol{A}[U-U_{1},V_{1}]$.
2.3.3 Relationship with Existing Recovery Guarantees
The sufficient conditions for recovery given by Theorem 2.1 and Theorem 2.2 can be shown to be extensions of several existing sufficient conditions for perfect recovery of dense subgraphs, planted cliques, and dense submatrices. Indeed, consider the $K=2$-case and set
$q:=p_{11},\hskip 18.06749ptp=p_{12}=p_{21}=p_{22}.$
In this case, the sufficient condition given in Theorem 2.1 immediately specializes to that given in (Bombina and Ames, 2020, Theorem 3.1). Moreover, placing further assumptions on $q$ and $p$, e.g., $q=1$, yield the recovery guarantees for the planted clique and maximum edge biclique problem proposed in Ames and Vavasis (2011). Simiilarly, the sufficient conditions for perfect recovery under the adversarially setting given in Theorem 2.2 directly specialize the sufficient conditions for perfect recovery of planted cliques and bicliques given in (Ames and Vavasis, 2011, Section 4.1 and Section 5.1).
3 Derivation of the Recovery Guarantees in the Random Case
The recovery guarantees established in this section are grounded in a rigorous duality-based analysis of the convex relaxation formulated in (3). In particular, we seek to characterize the conditions under which the solutions $\boldsymbol{\bar{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ not only satisfy feasibility but also exhibit optimality in the context of recovering the planted densest $(m_{1},n_{1})$-subgraph. This is achieved by constructing an explicit dual certificate that satisfies the Karush-Kuhn-Tucker (KKT) conditions for (3) given in the following theorem.
3.1 Optimality Conditions
Theorem 3.1.
The solutions $\boldsymbol{\bar{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ are the unique optimal solutions of (3), where $\boldsymbol{\bar{u}}$ and $\boldsymbol{\bar{v}}$ represent the characteristic vectors of $U_{1}$ and $V_{1}$, respectively, if and only if there exist dual multipliers $\lambda\geq 0$, $\boldsymbol{\Lambda}\in\mathbb{R}_{+}^{M\times N}$, $\boldsymbol{\Xi}\in\mathbb{R}_{+}^{M\times N}$, and $\boldsymbol{W}\in\mathbb{R}^{M\times N}$ satisfying the following conditions:
$\frac{\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}}{\sqrt{m_{1}n_{1}}}+\boldsymbol{W}-\lambda\boldsymbol{1}\boldsymbol{1}^{T}+\gamma\boldsymbol{1}\boldsymbol{1}^{T}-\boldsymbol{\Xi}+\boldsymbol{\Lambda}$ $=\boldsymbol{0}$ (8a)
$\mathrm{Tr}(\boldsymbol{\Lambda}^{T}(\boldsymbol{\bar{X}}-\boldsymbol{1}\boldsymbol{1}^{T}))$ $=0$ (8b)
$\mathrm{Tr}(\boldsymbol{\Xi}^{T}\boldsymbol{\bar{Y}})$ $=0$ (8c)
$\boldsymbol{W}^{T}\boldsymbol{\bar{u}}=\boldsymbol{0},\quad\boldsymbol{W}\boldsymbol{\bar{v}}=\boldsymbol{0},\quad\|\boldsymbol{W}\|$ $<1.$ (8d)
3.2 A Proposed Choice of Lagrange Multipliers
Theorem 3.1 presents a specialization of the Karush-Kuhn-Tucker (KKT) optimality conditions for the semidefinite program (3) (see Section 5.5.3 of Boyd (2004)). Our derivation of sufficient conditions for perfect recovery relies on solving the nonlinear system (8), which ensures both optimality and uniqueness of the solution given by Theorem 3.1. This solution corresponds to the planted submatrix with rows and columns indexed by $U_{1}$ and $V_{1}$, respectively. Let $\boldsymbol{\bar{u}}$ and $\boldsymbol{\bar{v}}$ denote the characteristic vectors of $\bar{U}_{1}$ and $\bar{V}_{1}$.
Under the assumptions of Theorem 2.1, we show that the solutions $\bar{\boldsymbol{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ are the unique optimal solutions of (3), and that the subgraph induced by the index sets $(m_{1},n_{1})$ represents the densest subgraph. Moreover, we demonstrate that $\{\bar{\boldsymbol{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})\}$ satisfy the sufficient conditions for optimality derived from the KKT framework.
The analysis is grounded in tools from convex optimization, spectral graph theory, and probabilistic analysis/sharp nonasymptotic bounds on the norm of random matrices with independent entries Bandeira and Van Handel (2016); Boucheron et al. (2013). By examining the primal-dual optimality conditions, we provide a concrete method to verify when a rank-one solution of the form $\boldsymbol{X}=\boldsymbol{u}\boldsymbol{v}^{T}$, where $\boldsymbol{u},\boldsymbol{v}$ corresponds to the characteristic vectors of the planted submatrix, satisfies the optimality criteria. The core idea is to carefully construct the dual variables to satisfy the complementary slackness and stationarity conditions of the KKT system while ensuring the feasibility of both primal and dual solutions.
This approach allows us to establish precise recovery guarantees for the combinatorial problem of submatrix localization under partial observation. By bridging the discrete structural properties of the graph and continuous convex relaxations, we provide a robust theoretical foundation for the proposed algorithmic framework under the probabilistic model.
3.2.1 Choice of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$
We can construct a solution of the KKT conditions (8) by constructing the matrices $\boldsymbol{W}$ and $\boldsymbol{\Xi}$ according to the following cases.
  • Case 1: If $(i,j)\in{\bar{U}}_{1}\times{\bar{V}}_{1}-\Omega$, then (8a) indicates that
    $W_{ij}$ $=\lambda-\gamma-\frac{1}{\sqrt{m_{1}n_{1}}}+\Xi_{ij}-\Lambda_{ij}$
    $=\lambda-\frac{1}{\sqrt{m_{1}n_{1}}}-\Lambda_{ij}$
    $=\bar{\lambda}-\Lambda_{ij},$
    where $\bar{\lambda}:=\lambda-\frac{1}{\sqrt{m_{1}n_{1}}}$, if we choose $\Xi_{ij}=\gamma.$
  • Case 2: If $i\in U_{1}$, $j\in V_{1}$ and $(i,j)\in\Omega$, then
    $X_{ij}=Y_{ij}=\frac{1}{\sqrt{m_{1}n_{1}}}.$
    The complementary slackness condition (8c) is only satisfied in this case if $\Xi_{ij}=0.$ In this case,
    $W_{ij}=\bar{\lambda}-\gamma-\Lambda_{ij}.$
  • Case 3: Suppose that $i\in U_{r},$ $j\in V_{s}$ for some $r,s\geq 2$. Then we choose
    $W_{ij}=\psi_{rs}\lambda\times\begin{cases}1,&\text{if }(i,j)\notin\Omega,\\ -\frac{p_{rs}}{1-p_{rs}},&\text{if }(i,j)\in\Omega,\end{cases}$ (9)
    where
    $\psi_{rs}=\begin{cases}1,&\text{if }p_{rs}\leq 1-p_{rs}\\ -\frac{1-p_{rs}}{p_{rs}},&\text{if }1-p_{rs}<p_{rs}.\end{cases}$ (10)
    We choose $\Xi_{ij}=W_{ij}-\lambda+\gamma$ according to (8a).
  • Case 4: Suppose that $i\in U_{1}$, $j\in V-V_{1}$.
    Suppose further that $p_{1s}\leq 1/2.$ In this case, we choose so $W_{ij}$ and $\Xi_{ij}$ as follows so that (8a) is satisfied:
    $W_{ij}$ $=\begin{cases}\lambda,&\text{if }(i,j)\notin\Omega,\\ -\lambda\left(\frac{\nu_{j}}{m_{1}-\nu_{j}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$ (11)
    $\Xi_{ij}$ $=\begin{cases}\gamma,&\text{if }(i,j)\notin\Omega,\\ \gamma-\lambda\left(\frac{m_{1}}{m_{1}-\nu_{j}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$ (12)
    where $m_{1}:=\lvert U_{1}\rvert$ and $\nu_{j}$ denotes the number of nonzero entries of the $j$th column of $\boldsymbol{A}$ in rows indexed by $U_{1}$.
    On the other hand, if $p_{1s}>1/2$, we choose
    $W_{ij}$ $=\begin{cases}\lambda\left(\frac{\nu_{j}-m_{1}}{\nu_{j}}\right),&\text{if }(i,j)\notin\Omega,\\ \lambda,&\text{if }(i,j)\in\Omega,\end{cases}$
    $\Xi_{ij}$ $=\begin{cases}\gamma-\lambda\left(\frac{m_{1}}{\nu_{j}}\right),&\text{if }(i,j)\notin\Omega,\\ \gamma,&\text{if }(i,j)\in\Omega.\end{cases}$
  • Case 5: Similarly, if $i\in U_{r}$, $r\geq 2$, $j\in V_{1}$ and $p_{r1}\leq 1/2$ we choose
    $W_{ij}$ $=\begin{cases}\lambda,&\text{if }(i,j)\notin\Omega,\\ -\lambda\left(\frac{\mu_{i}}{n_{1}-\mu_{i}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$
    $\Xi_{ij}$ $=\begin{cases}\gamma,&\text{if }(i,j)\notin\Omega,\\ \gamma-\lambda\left(\frac{n_{1}}{n_{1}-\mu_{i}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$
    where $n_{1}=\lvert V_{1}\rvert$ and $\mu_{i}$ denotes the number of nonzero entries of the $i$th row of $\boldsymbol{A}$ in columns indexed by $V_{1}$.
    If $p_{r1}>1/2$, we choose
    $W_{ij}$ $=\begin{cases}\lambda\left(\frac{\mu_{i}-n_{1}}{\mu_{i}}\right),&\text{if }(i,j)\notin\Omega,\\ \lambda,&\text{if }(i,j)\in\Omega.\end{cases}$
    $\Xi_{ij}$ $=\begin{cases}\gamma-\lambda\left(\frac{n_{1}}{\mu_{i}}\right),&\text{if }(i,j)\notin\Omega,\\ \gamma,&\text{if }(i,j)\in\Omega.\end{cases}$
This exhausts all possible values of $W_{ij}$ and $\Xi_{ij}$. We will use this choice to choose the remaining Lagrange multipliers.
3.2.2 Construction of $\boldsymbol{\Lambda}$
By construction, the proposed Lagrange multipliers satisfy the stationarity condition (8a) and the complementary slackness conditions (8b) and (8c). Moreover, for any $i\in U-U_{1}$, we have
$\boldsymbol{W}\boldsymbol{\bar{u}}]_{i}=\sum_{j\in V_{1}}W_{ij}=\mu_{i}\lambda-(n-\mu_{i})\lambda\left(\frac{\mu_{i}}{n_{1}-\mu_{i}}\right)=0.$
By a symmetric argument, we have $\boldsymbol{W}^{T}\boldsymbol{\bar{v}}]_{j}=0$ for all $j\in V-V_{1}$.
We next propose a choice of $\Lambda_{ij}$ for $i\in U_{1},$ $j\in V_{1}$ ensuring that $\boldsymbol{W}\boldsymbol{\bar{u}}]_{i}=0$ and $\boldsymbol{W}^{T}\boldsymbol{\bar{v}}]_{j}=0$. Verifying when these choices of $\boldsymbol{\Lambda}$ and $\boldsymbol{\Xi}$ are nonnegative and $\|\boldsymbol{W}\|<1$ with high probability will complete the proof.
We parametrize the (unknown) nonzero entries of $\boldsymbol{\Lambda}$ as
$\Lambda_{ij}=:y_{i}+z_{j}$ (13)
for each $(i,j)\in U_{1}\times V_{1}$. That is,
$\boldsymbol{\Lambda}(U_{1},V_{1})=\boldsymbol{y}\boldsymbol{1}^{T}+\boldsymbol{1}\boldsymbol{z}^{T}$
for some $\boldsymbol{y}\in\mathbf{R}^{m_{1}},$ $\boldsymbol{z}\in\mathbf{R}^{n_{1}}$. Under this parametrization of $\boldsymbol{\Lambda}$, the conditions $\boldsymbol{W}\boldsymbol{\bar{u}}=\boldsymbol{0}$ and $\boldsymbol{W}^{T}\boldsymbol{\bar{v}}=\boldsymbol{0}$ are satisfied if $\boldsymbol{y}$ and $\boldsymbol{z}$ are solutions of the linear system
$\left(\begin{array}[]{cc}n_{1}\boldsymbol{I}&\boldsymbol{1}\boldsymbol{1}^{T}\\ \boldsymbol{1}\boldsymbol{1}^{T}&m_{1}\boldsymbol{I}\end{array}\right)\left(\begin{array}[]{c}\boldsymbol{y}\\ \boldsymbol{z}\end{array}\right)=\left(\begin{array}[]{c}-\gamma\boldsymbol{\bar{\mu}}+n_{1}\bar{\lambda}\boldsymbol{1}\\ -\gamma\boldsymbol{\bar{\nu}}+m_{1}\bar{\lambda}\boldsymbol{1}\end{array}\right),$ (14)
where
$\bar{\mu}_{i}$ $:=n_{1}-\mu_{i},\hskip 25.29494pt\bar{\nu}_{j}:=m_{1}-\nu_{i}$
for all $i\in U_{1}$, and $j\in V_{1}$.
The coefficient matrix of this linear system is singular with null space spanned by $(\boldsymbol{1},-\boldsymbol{1})\in\mathbf{R}^{m_{1}+n_{1}}.$ However, the unique solution of the nonsingular linear system
$\left(\left(\begin{array}[]{cc}n_{1}\boldsymbol{I}&\boldsymbol{1}\boldsymbol{1}^{T}\\ \boldsymbol{1}\boldsymbol{1}^{T}&m_{1}\boldsymbol{I}\end{array}\right)+\theta\begin{pmatrix}\boldsymbol{1}\\ -\boldsymbol{1}\end{pmatrix}\begin{pmatrix}\boldsymbol{1}&-\boldsymbol{1}^{T}\end{pmatrix}\right)\left(\begin{array}[]{c}\boldsymbol{y}\\ \boldsymbol{z}\end{array}\right)=\left(\begin{array}[]{c}-\gamma\boldsymbol{\bar{\mu}}+n_{1}\bar{\lambda}\boldsymbol{1}\\ -\gamma\boldsymbol{\bar{\nu}}+m_{1}\bar{\lambda}\boldsymbol{1}\end{array}\right),$ (15)
obtained by perturbing (14) by $\theta(\boldsymbol{1},-\boldsymbol{1})$ is a solution of (14) for all $\theta>0$. Choosing $\theta=1$, the coefficient matrix of (15) becomes
$\left(\begin{array}[]{cc}n_{1}\boldsymbol{I}&\boldsymbol{1}\boldsymbol{1}^{T}\\ \boldsymbol{1}\boldsymbol{1}^{T}&m_{1}\boldsymbol{I}\end{array}\right)+\theta\begin{pmatrix}\boldsymbol{1}\\ -\boldsymbol{1}\end{pmatrix}\begin{pmatrix}\boldsymbol{1}&-\boldsymbol{1}^{T}\end{pmatrix}$ $=\left(\begin{array}[]{cc}n_{1}\boldsymbol{I}&\boldsymbol{1}\boldsymbol{1}^{T}\\ \boldsymbol{1}\boldsymbol{1}^{T}&m_{1}\boldsymbol{I}\end{array}\right)+\theta\left(\begin{array}[]{cc}\boldsymbol{1}\boldsymbol{1}^{T}&-\boldsymbol{1}\boldsymbol{1}^{T}\\ -\boldsymbol{1}\boldsymbol{1}^{T}&\boldsymbol{1}\boldsymbol{1}^{T}\end{array}\right)$
$=\left(\begin{array}[]{cc}n_{1}\boldsymbol{I}+\theta\boldsymbol{1}\boldsymbol{1}^{T}&\boldsymbol{1}\boldsymbol{1}^{T}-\theta\boldsymbol{1}\boldsymbol{1}^{T}\\ \boldsymbol{1}\boldsymbol{1}^{T}-\theta\boldsymbol{1}\boldsymbol{1}^{T}&m_{1}\boldsymbol{I}+\theta\boldsymbol{1}\boldsymbol{1}^{T}\end{array}\right)$
$=\left(\begin{array}[]{cc}n_{1}\boldsymbol{I}+\boldsymbol{1}\boldsymbol{1}^{T}&0\\ 0&m_{1}\boldsymbol{I}+\boldsymbol{1}\boldsymbol{1}^{T}\end{array}\right).$
In this case, (15) reduces to
$(n_{1}\boldsymbol{I}+\boldsymbol{1}\boldsymbol{1}^{T})\boldsymbol{y}$ $=-\gamma\boldsymbol{\bar{\mu}}+n_{1}\bar{\lambda}\boldsymbol{1}$
$(m_{1}\boldsymbol{I}+\boldsymbol{1}\boldsymbol{1}^{T})\boldsymbol{z}$ $=-\gamma\boldsymbol{\bar{\nu}}+m_{1}\bar{\lambda}\boldsymbol{1},$
which has unique solution given by
$\boldsymbol{y}$ $=(n_{1}\boldsymbol{I}+\boldsymbol{1}\boldsymbol{1}^{T})^{-1}(-\gamma\boldsymbol{\bar{\mu}}+n_{1}\bar{\lambda}\boldsymbol{1})$ (16)
$\boldsymbol{z}$ $=(m_{1}\boldsymbol{I}+\boldsymbol{1}\boldsymbol{1}^{T})^{-1}(-\gamma\boldsymbol{\bar{\nu}}+m_{1}\bar{\lambda}\boldsymbol{1}).$ (17)
Applying the Sherman-Morrison-Woodbury-Formula (Deng, 2011, Theorem 1.1) and (Golub and Van Loan, 2013, section 2.1.4) to Equation (16) yields
$\boldsymbol{y}$ $=\frac{1}{n_{1}}\left(\boldsymbol{I}-\frac{1}{m_{1}+n_{1}}\boldsymbol{1}\boldsymbol{1}^{T}\right)(-\gamma\boldsymbol{\bar{\mu}}+n_{1}\bar{\lambda}\boldsymbol{1})$
$=\frac{1}{n_{1}}\left(-\gamma\boldsymbol{\bar{\mu}}+\left(\frac{\gamma\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}}{m_{1}+n_{1}}\right)\boldsymbol{1}+\left(\frac{n_{1}^{2}}{m_{1}+n_{1}}\right)\bar{\lambda}\boldsymbol{1}\right).$ (18)
and
$\boldsymbol{z}=\frac{1}{m_{1}}\left(-\gamma\boldsymbol{\bar{\nu}}+\left(\frac{\gamma\boldsymbol{\bar{\nu}}^{T}\boldsymbol{1}}{m_{1}+n_{1}}\right)\boldsymbol{1}+\left(\frac{m_{1}^{2}}{m_{1}+n_{1}}\right)\bar{\lambda}\boldsymbol{1}\right).$ (19)
3.3 Nonnegativity of $\boldsymbol{\Lambda}$
The entries of $\boldsymbol{\bar{\mu}}$ and $\boldsymbol{\bar{\nu}}$ are binomial random variables corresponding to $n_{1}$ and $m_{1}$ independent Bernoulli trials, respectively, each with probability of success $1-p_{11}$. Thus,
$\mathbf{E}[\bar{\mu}_{i}]$ $=(1-p_{11})n_{1},$ $\operatorname{Var}[\bar{\mu}_{i}]$ $=p_{11}(1-p_{11})n_{1},$
$\mathbf{E}[\bar{\nu}_{i}]$ $=(1-p_{11})m_{1},$ $\operatorname{Var}[\bar{\nu}_{i}]$ $=p_{11}(1-p_{11})m_{1}.$
Moreover, the number of nonzero entries in the $U_{1}\times V_{1}$ block of $\boldsymbol{A}$ is equal to $\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}=\boldsymbol{\bar{\nu}}^{T}\boldsymbol{1}$. This is also binomially distributed, corresponding to $m_{1}n_{1}$ independent Bernoulli trials with probability of success $1-p_{11}$:
$\mathbf{E}[\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}]$ $=(1-p_{11})m_{1}n_{1},$ $\operatorname{Var}[\boldsymbol{\bar{\mu}}]_{i}$ $=p_{11}(1-p_{11})m_{1}n_{1}.$
Applying linearity of expectation implies that
$\mathbf{E}[y_{i}]$ $=\frac{n_{1}\bar{\lambda}}{m_{1}+n_{1}}+\frac{\gamma}{n_{1}(n_{1}+m_{1})}E[\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}]-\frac{\gamma}{n_{1}}E[\boldsymbol{\bar{\mu}}]_{i}$
$=\frac{n_{1}\bar{\lambda}e}{m_{1}+n_{1}}+\frac{\gamma m_{1}}{m_{1}+n_{1}}(1-p_{11})-\gamma(1-p_{11})$
$=\left(\frac{n_{1}}{m_{1}+n_{1}}\right)\left(\bar{\lambda}-\gamma(1-p_{11})\right)$ (20)
for all $i\in U_{1}$. Following an identical calculation, we obtain
$\mathbf{E}[z_{j}]=\frac{m_{1}}{m_{1}+n_{1}}(\bar{\lambda}-\gamma(1-p_{11}))$ (21)
for all $j\in V_{1}$. For any $\tau>0$, selecting
$\lambda=\frac{1}{\sqrt{m_{1}n_{1}}}+\gamma(1-p_{11})+\gamma\tau$ (22)
ensures that the entries of $\boldsymbol{\Lambda}$ are strictly positive in expectation.
Next, our focus shifts to determining $\tau$ in order to guarantee that both $\boldsymbol{y}$ and $\boldsymbol{z}$ have positive entries with high probability. To achieve this, we will repeatedly apply a the following specialization of the classical Bernstein inequality Boucheron et al. (2013)[Section 2.8] to bound the sum of independent Bernoulli random variables.
Lemma 3.1.
Consider a sequence of $k$ independent $\{0,1\}$ Bernoulli random variables, denoted by $x_{1},...,x_{k}$, each having a success probability of $\rho$. Let $s=\sum_{i=1}^{k}x_{i}$ represent the binomially distributed random variable indicating the number of successes. Then
$\operatorname{Pr}\left(\lvert s-\rho k\rvert>6\max\left\{\sqrt{\rho(1-\rho)k\log t},\log t\right\}\right)\leq 2t^{-6}$ (23)
for all $t>0$.
Applying Lemma 3.1 with $t=N^{*}=\max\{M,N\}$ and the union bound shows that
$\lvert{\bar{\mu}}_{i}-(1-p_{11})n_{1}\rvert$ $\leq 6\max\left\{\sqrt{\sigma_{11}^{2}n_{1}\log{N^{*}}},\log{N^{*}}\right\},$ (24)
$\lvert{\bar{\nu}}_{j}-(1-p_{11})m_{1}\rvert$ $\leq 6\max\left\{\sqrt{\sigma_{11}^{2}m_{1}\log{N^{*}}},\log{N^{*}}\right\},$ (25)
$\lvert\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}-m_{1}n_{1}(1-p_{11})\rvert$ $\leq 6\max\left\{\sqrt{\sigma_{11}^{2}m_{1}n_{1}\log N^{*}},\log{N^{*}}\right\}$ (26)
for all $i\in U_{1}$ and $j\in V_{1}$ with high probability. Combining (24), (25), and (26) and applying the triangle inequality shows that
[…]
$\leq 6\gamma\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{n_{1}}},\frac{\log{N^{*}}}{n_{1}}\right\}$
$\quad\quad+\frac{6\gamma}{m_{1}+n_{1}}\max\left\{\sqrt{\frac{\sigma_{11}^{2}m_{1}\log{N^{*}}}{n_{1}}},\frac{\log{N^{*}}}{n_{1}}\right\}$
$=\mathcal{O}\left(\gamma\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{n_{1}}},\frac{\log{N^{*}}}{n_{1}}\right\}\right)$ (27)
for all $i\in U_{1}$ w.h.p. By a symmetric argument
$\lvert z_{j}-\mathbf{E}[z_{j}]\rvert$ $=\mathcal{O}\left(\gamma\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{m_{1}}},\frac{\log{N^{*}}}{m_{1}}\right\}\right)$ (28)
for all $j\in V_{1}$ w.h.p. Substituting (27) and (28) into Equation (13) shows that
$\Lambda_{ij}$ $=y_{i}+z_{j}$
$\geq\mathbf{E}[y_{i}]-\lvert y_{i}-\mathbf{E}[y_{i}]\rvert+\mathbf{E}[z_{j}]-\lvert z_{j}-\mathbf{E}[z_{j}]\rvert$
$=\frac{n_{1}}{m_{1}+n_{1}}(\bar{\lambda}-\gamma(1-p_{11}))-\lvert y_{i}-\mathbf{E}[y_{i}]\rvert$
$\hskip 36.135pt+\frac{m_{1}}{m_{1}+n_{1}}(\bar{\lambda}-\gamma(1-p_{11}))-\lvert z_{j}-\mathbf{E}[z_{j}]\rvert$
$\geq\bar{\lambda}-\gamma(1-p_{11})-c_{\tau}\gamma\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{n_{1*}}},\frac{\log{N^{*}}}{n_{1*}}\right\}$
$=\gamma\tau-c_{\tau}\gamma\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{n_{1*}}},\frac{\log{N^{*}}}{n_{1*}}\right\}$
for some constant $c_{\tau}>0$ w.h.p. for all $i\in U_{1}$ and $j\in V_{1}$. Therefore, choosing sufficiently large $\tau>0$ ensures that $\boldsymbol{\Lambda}$ is nonnegative w.h.p. Specifically choosing
$\tau=c_{\tau}\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{n_{1*}}},\frac{\log{N^{*}}}{n_{1*}}\right\}$ (29)
guarantees that every entry of $\boldsymbol{\Lambda}$ is nonnegative w.h.p.
3.4 Nonnegativity of $\boldsymbol{\Xi}$
We next derive conditions on the regularization parameter $\gamma$ so that the entries of the dual variable $\boldsymbol{\Xi}$ are nonnegative with high probability.
It is essential to recall that ${\Xi}_{ij}$ takes values 0 and $\gamma$ for all $(i,j)$ except in Case 3 and the $(i,j)\in\Omega$ subcase in Cases 4 and 5 considered in the choice of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$ in Section 3.2.1. It remains to establish conditions under which $\Xi_{ij}\geq 0$ w.h.p. in these remaining cases.
3.4.1 Nonnegativity in Case 3
Consider $i\in U_{r}$ and $j\in V_{s}$ for $r,s\geq 2$. To determine sufficient conditions for nonnegativity of $\Xi_{ij}$, we must consider two cases dictating the choice of $\psi_{rs}$ and, hence, $W_{ij}$.
First, suppose that $p_{rs}\leq 1-p_{rs}$ or, equivalently, $p_{rs}\leq 1/2$. Recall that we choose $\psi=\psi_{rs}=1$ in this case. If $(i,j)\notin\Omega$, then
$\Xi_{ij}=W_{ij}-\lambda+\gamma=\lambda\psi-\lambda+\gamma=\gamma>0,$
by our choice of $W_{ij}$.
On the other hand, suppose that $(i,j)\in\Omega$. In this case,
$\Xi_{ij}$ $=W_{ij}-\lambda+\gamma=\gamma-\lambda\left(\psi\frac{p_{rs}}{1-p_{rs}}+1\right).$
Substituting $\lambda=\frac{1}{\sqrt{m_{1}n_{1}}}+\gamma(1-p_{11})+\gamma\tau$, we have
$(1-p_{rs})\Xi_{ij}$ $=\gamma(1-p_{rs})-\frac{1}{\sqrt{m_{1}n_{1}}}-\gamma(1-p_{11})-\gamma\tau$
$=\gamma(p_{11}-p_{rs}-\tau)-\frac{1}{\sqrt{m_{1}n_{1}}}.$
Under the assumption
$p_{11}-p_{rs}\geq 2\tau=2c_{\tau}\max\left\{\sqrt{\frac{\sigma_{11}^{2}\log{N^{*}}}{n_{1*}}},\frac{\log{N^{*}}}{n_{1*}}\right\},$ (30)
we have
$(1-p_{rs})\Xi_{ij}\geq\frac{\gamma}{2}(p_{11}-p_{rs})-\frac{1}{\sqrt{m_{1}n_{1}}}.$
This implies that $\Xi_{ij}\geq 0$ if
$\frac{\gamma}{2}(p_{11}-p_{rs})\geq\frac{1}{\sqrt{m_{1}n_{1}}}.$ (31)
We next consider the case when $p_{rs}>1/2$. In this case,
$\psi=\psi_{rs}=-\frac{1-p_{rs}}{p_{rs}}.$
If $(i,j)\notin\Omega$, we have
$\Xi_{ij}=W_{ij}-\lambda+\gamma=-\frac{\lambda}{p_{rs}}+\gamma.$
It follows that
$p_{rs}\Xi_{ij}$ $=-\lambda+p_{rs}\gamma$
$=p_{rs}\gamma-\frac{1}{\sqrt{m_{1}n_{1}}}-\gamma(1-p_{11})-\gamma\tau$
$>(1-p_{rs})\gamma-\frac{1}{\sqrt{m_{1}n_{1}}}-\gamma(1-p_{11})-\gamma\tau$
$\geq\frac{\gamma}{2}(p_{11}-p_{rs})-\frac{1}{\sqrt{m_{1}n_{1}}},$
under Assumption (30) since $p_{rs}>1-p_{rs}$. Thus, $\Xi_{ij}>0$ under Assumptions (30) and (31) in this case. Finally, if $(i,j)\in\Omega$ then
$\Xi_{ij}=-\lambda\left(\frac{p_{rs}}{1-p_{rs}}\right)\left(-\frac{1-p_{rs}}{p_{rs}}\right)-\lambda+\gamma=\gamma>0.$
3.4.2 Nonnegativity in Cases 4 and 5
We start with Case 4 in the construction of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$. Fix $i\in U_{1}$ and $j\in V_{s}$ for $s\geq 2$.
We first consider the case when $p_{1s}>1/2$ and $(i,j)\notin\Omega$. In this case,
$\Xi_{ij}$ $=\gamma-\lambda\frac{m_{1}}{\nu_{j}}.$ (32)
Note that $\Xi_{ij}$ is a monotonically increasing function of $\nu_{j}$ on the interval $0,m_{j})$. Therefore, $\Xi_{ij}$ is minimized when $\nu_{j}$ is minimized over this interval. Moreover, the Bernstein inequality implies that
$\nu_{j}$ $\geq p_{1s}m_{1}-6\max\left\{\sqrt{\sigma_{p_{1s}}^{2}m_{1}\log{N^{*}}},\log{N^{*}}\right\}.$ (33)
with high probability. Substituting (29) into (32) and rearranging we see that $\Xi_{ij}\geq 0$ w.h.p. if and only if $(\nu_{j}/m_{1})\Xi_{ij}\geq 0$, which holds if
$\frac{\nu_{j}}{m_{1}}\Xi_{ij}$ $\geq\frac{\gamma}{m_{1}}\left(p_{1s}m_{1}-6\max\left\{\sqrt{\sigma_{1s}^{2}m_{1}\log N^{*}},\log N^{*}\right\}\right)-\gamma\tau-\gamma(1-p_{11})-\frac{1}{\sqrt{m_{1}n_{1}}}$
$\geq\gamma\left(p_{11}-p_{1s}-\tau-6\max\left\{\sqrt{\sigma_{1s}^{2}m_{1}\log N^{*}},\log N^{*}\right\}\right)-\frac{1}{\sqrt{m_{1}n_{1}}}.$
If
$p_{11}-p_{1s}\geq 2\left(\tau+6\max\left\{\sqrt{\sigma_{1s}^{2}\frac{\log N^{*}}{m_{1}}},\frac{\log N^{*}}{m_{1}}\right\}\right),$ (34)
then we have $\Xi_{ij}\geq 0$ if
$\frac{\gamma}{2}(p_{11}-p_{1s})\geq\frac{1}{\sqrt{m_{1}n_{1}}}.$
Rearranging shows that $\Xi_{ij}\geq 0$ w.h.p. if (34) holds and we choose
$\gamma\geq\frac{2}{(p_{11}-p_{1s})\sqrt{m_{1}n_{1}}}.$ (35)
Establishing nonnegativity of $\boldsymbol{\Xi}$ under Case 5 in Section 3.2.1 follows a similar argument. Suppose that $i\in U_{1}$, $j\in V_{r}$ for $r\geq 2$ such that $(i,j)\in\Omega$. Then $\Xi_{ij}\geq 0$ with high probability if
$p_{11}-p_{r1}\geq 2\left(\tau+6\max\left\{\sqrt{\sigma_{1s}^{2}\frac{\log N^{*}}{n_{1}}},\frac{\log N^{*}}{n_{1}}\right\}\right)$ (36)
and
$\gamma\geq\frac{2}{(p_{11}-p_{r1})\sqrt{m_{1}n_{1}}}.$ (37)
This establishes that $\boldsymbol{\Xi}$ has nonnegative entries in all cases with high probability.
3.5 A Bound on Matrix $\boldsymbol{W}$
We conclude the proof by establishing a sufficient condition involving the parameters $m_{1},n_{1},M,N,p_{rs}$ to guarantee that the constructed $\boldsymbol{W}$, as described above, satisfies the condition $\lVert\boldsymbol{W}\rVert\leq 1$ with high probability.
3.5.1 A Decomposition of $\boldsymbol{W}$
We establish an upper bound on $\lVert\boldsymbol{W}\rVert$ by utilizing the decomposition $\boldsymbol{W}=\lambda\boldsymbol{R}+\lambda\boldsymbol{S}$ with matrices $\boldsymbol{R}$ and $\boldsymbol{S}$ constructed as follows.
  • For each $i\in U_{1}$, $j\in V_{1}$, we choose $R_{ij}=\frac{W_{ij}}{\gamma}$ and $S_{ij}=0$
  • We set $R_{ij}=0$ $S_{ij}=\frac{W_{ij}}{\lambda}$ for all $(i,j)\notin U_{1}\times V_{1}$.
We will bound $\lVert{\boldsymbol{W}}\rVert$ by obtaining bounds on each of $\lVert{\boldsymbol{R}}\rVert$ and $\lVert{\boldsymbol{S}}\rVert$ and then applying the triangle inequality. To do so, we will make repeated use of the following bound on the following matrix concentration inequality, which provides a bound on the spectral norm of random matrices with mean-zero entries.
Lemma 3.2.
Let $\boldsymbol{A}=[a_{ij}]\in\mathbf{R}^{m\times n}$ be a random matrix with entries satisfying $\mathbf{E}[a_{ij}]=0$, $\operatorname{Var}[a_{ij}]\leq\sigma^{2}$, and $\lvert a_{ij}\rvert\leq B$ for all $i=1,\dots,m$ and $j=1,\dots,n$ for some positive constants $\sigma^{2}$ and $B$. Let
$\sigma_{1}$ $:=\max_{i}\sqrt{\sum_{j}\mathbf{E}[a_{ij}^{2}]}\leq\sigma\sqrt{n}$
$\sigma_{2}$ $:=\max_{j}\sqrt{\sum_{i}\mathbf{E}[a_{ij}^{2}]}\leq\sigma\sqrt{m}$
There exist a constant $C>0$ such that
$\Pr\left(\lVert{\boldsymbol{A}}\rVert\geq C\max\left\{\sigma_{1}+\sigma_{2},\sqrt{B\log{t}}\right\}\right)\leq\max\{n_{r},m_{s}\}t^{-7}$
for all $t>0$.
Lemma 3.2 is a special case of the matrix concentration inequality given in (Van Handel and Youssef, 2017, Remark 4.11). Note there is no assumption of symmetry, i.e., that $a_{ij}=a_{ji}$, in the hypothesis of Lemma 3.2.
3.5.2 A Bound on $\boldsymbol{R}$
The following lemma gives a bound on $\lVert{\boldsymbol{R}}\rVert$.
Lemma 3.3.
If the matrix $\boldsymbol{W}$ is constructed based on Cases 1 through 6 for a matrix $\boldsymbol{A}$ sampled from the planted dense submatrix model, then there exists a constant $c_{R}$ such that
$\|\boldsymbol{R}\|\leq c_{R}\max\left\{\sqrt{\sigma_{11}^{2}n_{1}^{*}\log N},\sqrt{\frac{n_{1}^{*}}{n_{1*}}}\log N\right\}$ (38)
with high probability, where $N^{*}=\max\{M,N\},$ $n_{1}^{*}=\max\{m_{1},n_{1}\}$, and $n_{1*}=\min\{m_{1},n_{1}\}.$
Proof.
By construction, $\|\boldsymbol{R}\|=\|\boldsymbol{R}(U_{1},V_{1})\|$. Note that
$\boldsymbol{R}(U_{1},V_{1})=\frac{1}{n_{1}}\boldsymbol{\bar{\mu}}\boldsymbol{1}^{T}+\frac{1}{m_{1}}\boldsymbol{1}\boldsymbol{\bar{\nu}}^{T}-\frac{\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}}{m_{1}n_{1}}\boldsymbol{1}\boldsymbol{1}^{T}-\boldsymbol{H}$
by our earlier formulas for $\boldsymbol{y}$ and $\boldsymbol{z}$, where $\boldsymbol{H}$ is defined by $H_{ij}=1$ if $(i,j)\in\Omega$ and $H_{ij}=0$ otherwise. This suggests that we can decompose $\boldsymbol{R}(U_{1},V_{1})$ as $\boldsymbol{R}(U_{1},V_{1})=\boldsymbol{Q}_{1}+\boldsymbol{Q}_{2}+\boldsymbol{Q}_{3}+\boldsymbol{Q}_{4}$, where
$\boldsymbol{Q}_{1}$ $=(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T}-\boldsymbol{H}$
$\boldsymbol{Q}_{2}$ $=\frac{1}{n_{1}}\boldsymbol{\bar{\mu}}\boldsymbol{1}^{T}-(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T}$
$\boldsymbol{Q}_{3}$ $=\frac{1}{m_{1}}\boldsymbol{1}\boldsymbol{\bar{\nu}}^{T}-(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T}$
$\boldsymbol{Q}_{4}$ $=\left(\frac{(1-p_{11})m_{1}n_{1}-\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}}{m_{1}n_{1}}\right)\boldsymbol{1}\boldsymbol{1}^{T}$
We start by bounding $\boldsymbol{Q}_{1}$. Notice that $\boldsymbol{Q}_{1}=(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T}-\boldsymbol{H}$ is centered with mean-zero entries and variance $\sigma_{{11}}^{2}=p_{11}(1-p_{11})$. Applying Lemma 3.2 with $B=1$ and $t=N^{*}=\max\{N,M\}$, we have
$\lVert\boldsymbol{Q}_{1}\rVert$ $=\lVert(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T}-\boldsymbol{H}\rVert=\mathcal{O}\left(\max\left\{\left(\sqrt{n_{1}}+\sqrt{m_{1}}\right)\sigma_{{11}},\sqrt{\log{N^{*}}}\right\}\right)$
$=\mathcal{O}\left(\max\left\{\sqrt{n_{1}^{*}\sigma^{2}_{{11}}},\sqrt{\log{N^{*}}}\right\}\right)$ (39)
Next, taking $\boldsymbol{Q}_{2}=\frac{1}{n_{1}}(\boldsymbol{\bar{\mu}}\boldsymbol{1}^{T}-(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T}$ and applying Lemma 3.1, we have
$\lVert\boldsymbol{Q}_{2}\rVert^{2}$ $=\frac{1}{n_{1}}\|\boldsymbol{\bar{\mu}}-(1-p_{11})n_{1}\|\|\boldsymbol{1}\|$
$=\frac{1}{\sqrt{n_{1}}}\|\boldsymbol{\bar{\mu}}-(1-p_{11})n_{1}\|$
$=\mathcal{O}\left(\sqrt{m_{1}}\cdot\frac{1}{\sqrt{n_{1}}}\max\left\{\sqrt{\sigma^{2}_{11}n_{1}\log N},\log N\right\}\right)$
$=\mathcal{O}\left(\max\left\{\sqrt{\sigma^{2}_{11}m_{1}\log N},\sqrt{\frac{m_{1}}{n_{1}}}\log N\right\}\right)$ (40)
since
$\lvert\boldsymbol{\bar{\mu}}_{i}-(1-p_{11})n_{1}\rvert=\mathcal{O}\left(\max\left\{\sqrt{\sigma^{2}_{11}n_{1}\log N},\log N\right\}\right)$
w.h.p. for all $i\in U_{1}$ by Bernstein’s inequality.
By an identical argument $\boldsymbol{Q}_{3}=\frac{1}{m_{1}}(\boldsymbol{1}\mathbf{\boldsymbol{\bar{\nu}}^{T}}-(1-p_{11})\boldsymbol{1}\boldsymbol{1}^{T})$ satisfies
$\|\boldsymbol{Q}_{3}\|$ $=\mathcal{O}\left(\max\left\{\sqrt{\sigma^{2}_{11}m_{1}\log N},\log N\right\}\right)$ (41)
with high probability It remains to bound $\|\boldsymbol{Q}_{4}\|$. We first note that
$\|\boldsymbol{Q}_{4}\|=\sqrt{m_{1}n_{1}}\left\lvert\frac{(1-p_{11})m_{1}n_{1}-\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}}{m_{1}n_{1}}\right\rvert$.
Moreover, $\boldsymbol{\bar{\mu}}^{T}\boldsymbol{1}$ is a binomially distributed random variable corresponding to $m_{1}n_{1}$ independent Bernoulli trials with probability of success $1-p_{11}$. Applying Lemma 3.1 one more time shows that
$\lVert\boldsymbol{Q}_{4}\rVert$ $=\mathcal{O}\left(\max\left\{\sqrt{\sigma_{11}^{2}\log{N}},\frac{\log N}{\sqrt{m_{1}n_{1}}}\right\}\right)$ (42)
with high probability.
Applying the triangle inequality, (39), (40), (41), (42), and the union bound shows that
$\lVert\boldsymbol{R}\rVert$ $=\mathcal{O}\left(\max\left\{\sqrt{\sigma_{11}^{2}n_{1}^{*}\log N},\sqrt{\frac{n_{1}^{*}}{n_{1*}}}\log N\right\}\right).$ (43)
with high probability. ∎
3.5.3 A Bound on $\boldsymbol{S}$
We have the following bound on the second matrix, $\boldsymbol{S}$, in the decomposition of $\boldsymbol{W}$.
Lemma 3.4.
If the matrix $\boldsymbol{W}$ is constructed according to Cases 1 through 6 for a matrix $\boldsymbol{A}$ sampled from the planted dense submatrix model, then there exist constant $c_{S}$ such that
$\|\boldsymbol{S}\|\leq c_{S}\max\left\{\sqrt{\tilde{\sigma}^{2}_{*}N^{*}\log N^{*}},(\log N^{*})^{3/2}\right\}$ (44)
with high probability.
The rest of this section consists of a proof of Lemma 3.4. To establish the desired bound on $\boldsymbol{S}$, we initially approximate $\boldsymbol{S}$ with a random matrix having mean-zero entries. Specifically, we define $\boldsymbol{S}_{1}$ as the random matrix constructed in the following manner:
  • For each $i\in U_{r}$, $j\in V_{s}$ such that $r,s\geq 2$, we let
    $\boldsymbol{S}_{1}]_{ij}=S_{ij}=\frac{W_{ij}}{\lambda}=\psi_{rs}\times\begin{cases}1,&\text{if }(i,j)\notin\Omega,\\ -\frac{p_{rs}}{1-p_{rs}},&\text{if }(i,j)\in\Omega,\end{cases}$
    where $\psi_{rs}$ is defined as in (10).
  • For $i\in U_{1}$, $j\in V_{s}$, $s\geq 2$, we independently sample $\boldsymbol{S}_{1}]_{ij}$ from the Bernoulli distribution such that
    $\boldsymbol{S}_{1}]_{ij}=\psi_{1s}\times\begin{cases}1,&\text{with probability $p_{1s}$},\\ -\frac{p_{1s}}{1-p_{1s}},&\text{with probability }1-p_{1s},\end{cases}$
    where
    $\psi_{1s}=\begin{cases}1,&\text{if }p_{1s}\leq 1-p_{1s}\\ -\frac{1-p_{1s}}{p_{1s}},&\text{if }1-p_{1s}<p_{1s}.\end{cases}$
  • For $i\in U_{r}$, $j\in V_{1}$, $r\geq 2$, we sample $\boldsymbol{S}_{1}]_{ij}$ from the Bernoulli distribution such that
    $\boldsymbol{S}_{1}]_{ij}=\psi_{r1}\times\begin{cases}1,&\text{with probability $p_{r1}$},\\ -\frac{p_{r1}}{1-p_{r1}},&\text{with probability }1-p_{r1},\end{cases}$
    where
    $\psi_{r1}=\begin{cases}1,&\text{if }p_{r1}\leq 1-p_{r1}\\ -\frac{1-p_{r1}}{p_{r1}},&\text{if }1-p_{r1}<p_{r1}.\end{cases}$
Note that entries in the $(U_{r},V_{s})$-block are random variables with variance
$\tilde{\sigma}_{rs}^{2}:=\min\left\{\frac{p_{rs}}{1-p_{rs}},\frac{1-p_{rs}}{p_{rs}}\right\}$
for all $(r,s)\neq(1,1).$ By construction, the entries of $\boldsymbol{S}_{1}$ are mean-zero independent random variables, each having variance bounded above by $\tilde{\sigma}^{2}$, where
$\tilde{\sigma}^{2}=\max_{\begin{subarray}{c}r,s\geq 2\\ r\neq s\end{subarray}}\left\{\tilde{\sigma}^{2}_{rs}\right\}$
Applying Lemma 3.2 with $B=\max\{1,\tilde{\sigma}^{2}\}\leq 1$ and $t=N^{*}=\max\{N,M\}$, shows that
$\lVert\boldsymbol{S}_{1}\rVert=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}},\sqrt{\log N^{*}}\right\}\right)$ (45)
with high probability.
The rest of the proof demonstrates that $\boldsymbol{S}$ is well approximated by $\boldsymbol{S}_{1}$. That is, we finish the proof by bounding $\lVert\boldsymbol{S}-\boldsymbol{S}_{1}\rVert$. To do so, we consider the matrices $\boldsymbol{S}_{2}$ and $\boldsymbol{S}_{3}$ defined as follows:
$[\boldsymbol{S}_{2}]_{ij}$ $=\begin{cases}[\boldsymbol{S}-\boldsymbol{S}_{1}]_{ij},&\text{if }i\in U_{1},j\in V_{s},s\geq 2,\\ 0,&\text{otherwise}.\end{cases}$
$[\boldsymbol{S}_{3}]_{ij}$ $=\begin{cases}[\boldsymbol{S}-\boldsymbol{S}_{1}]_{ij},&\text{if }i\in U_{r},j\in V_{1},r\geq 2,\\ 0,&\text{otherwise}.\end{cases}$
To bound the norm of $\boldsymbol{S}_{2}$ and the norm of $\boldsymbol{S}_{3}$, we will apply the following lemma, which offers bounds on the spectral norm of random matrices with this particular form. A proof of Lemma 3.5 is given in Appendix A
Lemma 3.5.
Let $V_{1},V_{2},\dots,V_{K}$ be a partition of $[N]$. Let $\boldsymbol{\Theta}$ be a random $n\times N$ matrix such that
$\Theta_{ij}=\begin{cases}1,&\text{with probability }p_{s}\\ -\frac{p_{s}}{1-p_{s}},&\text{with probability }1-p_{s}\end{cases}$ (46)
for all $i\in[n]$, $j\in V_{s}$ for probability $p_{s}\in[0,1/2]$. Let $\tilde{\boldsymbol{\Theta}}$ be the random matrix defined by
[…]
for each $i\in[n]$ and $j\in V_{s}$, where $n_{j}$ is the number of $1$s in the $j$th column of $\boldsymbol{\Theta}$. Then there is constant $c>0$ such that
$\operatorname{Pr}\left(\lVert\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}\rVert\geq c\max\left\{\sqrt{\hat{\sigma}^{2}N\log N},(\log N)^{3/2}\right\}\right)=\mathcal{O}\left(N^{-5}\right),$
where
$\hat{\sigma}^{2}:=\max_{s}\frac{p_{s}}{1-p_{s}}.$
Note that $\boldsymbol{S}_{2}$ has the same form as $\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}$ in Lemma 3.5. Indeed, if $p_{1s}\leq 1/2$ then $S_{ij}$ is sampled according to the distribution defining $\boldsymbol{\Theta}$ given in (46) and $\boldsymbol{S}_{1}]_{ij}$ is sampled according to (47) with respect to $p_{s}=p_{1s}$. On the other hand, if $p_{1s}>1/2$ then $S_{ij}$ and $\boldsymbol{S}_{1}]_{ij}$ are sampled according to (46) and (47), respectively, with $p_{s}=1-p_{1s}<1/2$. Thus, $\boldsymbol{S}_{2}$ is of the form $\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}$ in Lemma 3.5 with $n=n_{1}$, $N=N-n_{1}$. It follows immediately that
$\|\boldsymbol{S}_{2}\|$ $=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}(N-n_{1})\log(N-n_{1})},(\log(N-n_{1}))^{3/2}\right\}\right)$
$=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}\log N^{*}},(\log N^{*})^{3/2}\right\}\right).$ (48)
with high probability. By an identical calculation,
$\|\boldsymbol{S}_{3}\|$ $=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}(M-m_{1})\log(M-m_{1})},(\log(M-m_{1}))^{3/2}\right\}\right)$
$=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}\log N^{*}},(\log N^{*})^{3/2}\right\}\right)$ (49)
with high probability. Combining (45), (48), and (49) shows that
$\|\boldsymbol{S}\|=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}_{*}N^{*}\log N^{*}},(\log N^{*})^{3/2}\right\}\right)$
with high probability.
3.5.4 Completing the Bound on $\boldsymbol{W}$
Without loss of generality, we assume $N=N^{*}$ and $m_{1}\leq n_{1}$for notational simplicity. Applying the triangle inequality with the bounds given by Lemma 3.3 and Lemma 3.4 shows that
$\|\boldsymbol{W}\|$ $\leq\gamma\|\boldsymbol{R}\|+\lambda\|\boldsymbol{S}\|$
$\leq\gamma~c_{R}\max\left\{\max_{j}\sqrt{\tilde{\sigma}_{jj}^{2}n_{j}^{*}\log N},\sqrt{\frac{n_{1}}{m_{1}}}\log N\right\}$
$\hskip 18.06749pt+\lambda~c_{S}\max\left\{\sqrt{\tilde{\sigma}_{*}^{2}N\log N},(\log N)^{3/2}\right\}$ (50)
with high probability. Recall that
$\lambda$ $=\frac{1}{\sqrt{m_{1}n_{1}}}+\gamma(1-p_{11})+\gamma\tau$
$=\frac{1}{\sqrt{m_{1}n_{1}}}+\mathcal{O}\left(\gamma\max\left\{(1-p_{11}),\frac{\sigma_{11}^{2}\log N}{n_{\min}},\frac{\log N}{n_{\min}}\right\}\right)$
$=\frac{1}{\sqrt{m_{1}n_{1}}}+\mathcal{O}(\gamma)$
by our choice of $\tau$ and the assumption that $(\log N)^{3/2}=\mathcal{O}(n_{\min})$.
It follows that $\|\boldsymbol{W}\|\leq 1$ if we choose $\gamma\leq c_{\gamma}/\sqrt{m_{1}n_{1}}$ for sufficiently small $c_{\gamma}>0$ if
$\sqrt{\tilde{\sigma}_{*}^{2}N\log N}=\mathcal{O}(\sqrt{m_{1}n_{1}}).$ (51)
Indeed, (51) implies that
$\frac{1}{\sqrt{m_{1}n_{1}}}\sqrt{\tilde{\sigma}_{jj}^{2}n_{j}^{*}\log N}\leq\frac{1}{\sqrt{m_{1}n_{1}}}\sqrt{\tilde{\sigma}_{*}^{2}N\log N}=\mathcal{O}(1)$
for all $j$, since $n_{j}^{*}\leq N$ and $\tilde{\sigma}_{jj}^{2}\leq\tilde{\sigma}_{*}^{2}$. Moreover,
$\frac{1}{\sqrt{m_{1}n_{1}}}\sqrt{\frac{n_{1}}{m_{1}}}\log N=\frac{\log N}{m_{1}}=\mathcal{O}(1).$
Substituting these bounds in (50) shows that
$\|\boldsymbol{W}\|=\mathcal{O}(c_{\gamma})$
w.h.p. and, hence, choosing sufficiently small $c_{\gamma}$ ensures that $\|\boldsymbol{W}\|<1$ w.h.p.
4 Proof of Recovery in the Adversarial Case
The proof the sufficient condition in the adversarial case follows a similar structure that of Section 3. The proof relies on verifying that a particular choice of Lagrange multipliers satisfies the uniqueness and optimality conditions given in Theorem 3.1. As in Section 3, we start with a choice of $\boldsymbol{W}$.
  • Case 1: If $(i,j)\in{\bar{U}}_{1}\times{\bar{V}}_{1}-\Omega$, then (8a) indicates that
    $\Xi_{ij}$ $=\gamma$ $W_{ij}$ $=\bar{\lambda}-\Lambda_{ij}.$
  • Case 2: If $(i,j)\in\Omega\cap({\bar{U}}_{1}\times{\bar{V}}_{1})$, then
    $\Xi_{ij}$ $=0$ $W_{ij}$ $=\bar{\lambda}-\gamma-\Lambda_{ij}.$
  • Case 3: Suppose that $i\in U_{q},$ $j\in V_{s}$ for some $q,s\geq 2$. Then we choose
    $W_{ij}$ $=\begin{cases}\lambda,&\text{if }(i,j)\notin\Omega,\\ 0,&\text{if }(i,j)\in\Omega,\end{cases}$ $\Xi_{ij}$ $=\begin{cases}\gamma,&\text{if }(i,j)\notin\Omega,\\ \gamma-\lambda,&\text{if }(i,j)\in\Omega\end{cases}$
    according to (8a).
  • Case 4: Suppose that $i\in U_{1}$, $j\in V-V_{1}$.
    $W_{ij}$ $=\begin{cases}\lambda,&\text{if }(i,j)\notin\Omega,\\ -\lambda\left(\frac{\nu_{j}}{m_{1}-\nu_{j}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$ $\Xi_{ij}$ $=\begin{cases}\gamma,&\text{if }(i,j)\notin\Omega,\\ \gamma-\lambda\left(\frac{m_{1}}{m_{1}-\nu_{j}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$
    where $m_{1}:=\lvert U_{1}\rvert$ and $\nu_{j}$ denotes the number of nonzero entries of the $j$th column of $\boldsymbol{A}$ in rows indexed by $U_{1}$.
  • Case 5: Similarly, if $i\in U-U_{1}$, $j\in V_{1}$, we choose
    $W_{ij}$ $=\begin{cases}\lambda,&\text{if }(i,j)\notin\Omega,\\ -\lambda\left(\frac{\mu_{i}}{n_{1}-\mu_{i}}\right),&\text{if }(i,j)\in\Omega,\end{cases}$ $\Xi_{ij}$ $=\begin{cases}\gamma,&\text{if }(i,j)\notin\Omega,\\ \gamma-\lambda\left(\frac{n_{1}}{n_{1}-\mu_{i}}\right),&\text{if }(i,j)\in\Omega.\end{cases}$
4.1 Nonnegativity of $\boldsymbol{\Lambda}$ in the Adversarial Case
As in Section 3.2.2, we choose
$\Lambda_{ij}=y_{i}+z_{j}=\bar{\lambda}-\gamma\bigg(\frac{\bar{\mu}_{i}}{n_{1}}+\frac{\bar{\nu}_{j}}{m_{1}}-\frac{\bar{r}_{11}}{m_{1}n_{1}}\bigg)\\$
for all $(i,j)\in\bar{U}_{1}\times\bar{V}_{1}$ so that $\boldsymbol{W}^{T}\boldsymbol{\bar{u}}=\boldsymbol{0}$ and $\boldsymbol{W}\boldsymbol{\bar{v}}=\boldsymbol{0}$, where $r_{11}$ denotes the number of 0 entries in the $(\bar{U}_{1},\bar{V}_{1})$-block of $\boldsymbol{A}$.
By assumption, we have $\bar{\mu}_{i}\geq(1-\tilde{\delta})n_{1}$ and $\bar{\nu}_{j}\geq(1-\tilde{\delta})m_{1}$ for all $(i,j)$. It follows that
$\Lambda_{ij}$ $\geq\bar{\lambda}-2(1-\tilde{\delta})\gamma\geq 0$
if
$\lambda\geq\frac{1}{m_{1}n_{1}}+2(1-\tilde{\delta})\gamma.$ (52)
4.2 Nonnegativity of $\boldsymbol{\Xi}$ in the Adversarial Case
As in Section 3.4, we need only consider Cases 3,4, and 5 in the construction of $\boldsymbol{\Xi}$.
We begin with Case 3. Suppose that $(i,j)\in\Omega\cap U_{q}\times V_{s}$. In this case, we have
$\Xi_{ij}=\gamma-\lambda=\geq 0$
if $\gamma\geq 0$. Enforcing (52) shows that it suffices to choose $\gamma$ satisfying
$(2\tilde{\delta}-1)\gamma\geq\frac{1}{\sqrt{m_{1}n_{1}}}.$ (53)
Note that (53) is only satisfiable if $\tilde{\delta}>1/2$.
We next consider Case 4, i.e., when $i\in U_{1}$, $j\in V-V_{1}$ and $(i,j)\in\Omega$. In this case,
$\Xi_{ij}=\gamma-\lambda\left(\frac{m_{1}}{m_{1}-\nu_{j}}\right)\geq\gamma-\frac{\lambda}{1-\delta}$
by the assumption that $\nu_{j}\leq\delta m_{1}$. Again, enforcing (52) shows that $\Xi_{ij}$ in this case if
$(2\tilde{\delta}-\delta-1)\gamma\geq\frac{1}{\sqrt{m_{1}n_{1}}}.$ (54)
Finally, a symmetric argument shows that $\Xi_{ij}\geq 0$ if (54) is satisfied when $i\in U-U_{1}$, $j\in V_{1}$, and $(i,j)\in\Omega$. Choosing $\gamma$ according to (7) ensures that (53) and (54) are satisfied. Moreover, we choose
$\lambda=\frac{1}{m_{1}n_{1}}+\frac{2(1-\tilde{\delta})}{2\tilde{\delta}-\delta-1}\cdot\frac{1}{\sqrt{m_{1}n_{1}}}:=\frac{1}{m_{1}n_{1}}+\frac{\tilde{c}(1-\tilde{\delta})}{\sqrt{m_{1}n_{1}}},$
where $\tilde{c}:=2/(2\tilde{\delta}-\delta-1)$, which satisfies (52).
4.3 A Bound on $\boldsymbol{W}$ in the Adversarial Case
It remains to establish conditions when $\|\boldsymbol{W}\|\leq 1$ under the assumptions given in Section 2.3. To do so, we decompose $\boldsymbol{W}$ as $\boldsymbol{W}=\boldsymbol{W}_{1}+\boldsymbol{W}_{2}+\boldsymbol{W}_{3}$, where
$\boldsymbol{W}_{1}[U_{q},V_{s}]$ $=\begin{cases}\boldsymbol{W}[U_{q},V_{q}],&\text{if }~q=s,\\ 0,&o/w,\end{cases}$
$\boldsymbol{W}_{2}[U_{q},V_{s}]$ $=\begin{cases}\boldsymbol{W}[U_{1},V_{s}],&\text{if }~q=1,s\geq 2\\ \boldsymbol{W}[U_{q},V_{1}],&\text{if }~q\geq 2,s=1\\ 0,&o/w,\end{cases}$
$\boldsymbol{W}_{3}[U_{q},V_{s}]$ $=\begin{cases}\boldsymbol{W}[U_{q},V_{s}],&\text{if }~q,s\geq 2,q\neq s\\ 0,&o/w,\end{cases}$
We will bound each of $\|\boldsymbol{W_{1}}\|,\|\boldsymbol{W_{2}}\|,$ and $\|\boldsymbol{W_{3}}\|$ individually and use the triangle inequality to bound $\|\boldsymbol{W}\|$. We start with $\|\boldsymbol{W_{2}}\|$ and the following initial bound:
$\|\boldsymbol{W}_{2}\|^{2}\leq\|\boldsymbol{W}_{2}\|^{2}_{F}=\sum^{k}_{s=2}\left(\|\boldsymbol{W}[U_{1},V_{s}]\|^{2}_{F}+\|\boldsymbol{W}[U_{s},V_{1}]\|^{2}_{F}\right).$
The first summand is bounded above by
$\sum^{k}_{s=2}\|\boldsymbol{W}[U_{1},V_{s}]\|^{2}_{F}$ $=\sum^{k}_{i\in U_{1}}\sum^{k}_{j\in V-V_{1}}W_{ij}^{2}$
$=\lambda^{2}\sum^{k}_{j\in V-V_{1}}\bigg(v_{j}+(m_{1}-v_{j})\frac{v_{j}^{2}}{(m_{1}-v_{j})^{2}}\bigg)$
$=\lambda^{2}\sum^{k}_{j\in V-V_{1}}v_{j}\bigg(1+\frac{v_{j}}{m_{1}-v_{j}}\bigg)$
$\leq\lambda^{2}\sum^{k}_{j\in V-V_{1}}\frac{v_{j}}{1-\delta}\leq\lambda^{2}\frac{{r}_{1}}{1-\delta}.$
By an identical argument, we have
$\sum^{k}_{s=2}\|\boldsymbol{W}[U_{s},V_{1}]\|^{2}_{F}\leq\lambda^{2}\frac{{r}_{2}}{1-\delta}.$
Combining these two bounds, we have
$\|\boldsymbol{W_{2}}\|^{2}\leq\frac{\lambda^{2}}{1-\delta}\left(r_{1}+r_{2}\right).$ (55)
We bound $\boldsymbol{W_{3}}$ using a similar argument as
$\|\boldsymbol{W}_{3}\|^{2}\leq\|\boldsymbol{W}_{3}\|^{2}_{F}\leq\lambda^{2}{r}_{3}.$ (56)
It remains to bound $\boldsymbol{W_{1}}$.
To do so, we note that $\boldsymbol{W_{1}}$ is block diagonal and, therefore,
$\|\boldsymbol{W}_{1}\|=\max_{q=1,\dots,k}\|\boldsymbol{W}[U_{q},V_{q}]\|.$
For $q\geq 2$, there are at most $r_{qq}$ entries of $\boldsymbol{W}[U_{q},V_{q}]$ equal to $\lambda$, while the remaining entries are zeros. Thus,
$\|\boldsymbol{W}[U_{q},V_{q}]\|^{2}\leq\|\boldsymbol{W}[U_{q},V_{q}]\|^{2}_{F}\leq\lambda^{2}r_{qq}.$ (57)
On the other hand,
$\boldsymbol{W}(u_{1},v_{1})=\bar{\lambda}\boldsymbol{1}\boldsymbol{1}-\boldsymbol{1}\boldsymbol{y}^{T}+\boldsymbol{1}\boldsymbol{z}^{T}-\gamma\bar{\boldsymbol{A}}(u_{1},v_{1}).$
We decompose $\boldsymbol{W_{1}}$ further as $\boldsymbol{W_{1}}[U_{1},V_{1}]=\boldsymbol{\tilde{W}_{1}}+\boldsymbol{\tilde{W}_{2}}$, where
$\tilde{\boldsymbol{W}}_{1}$ $=\gamma\left(\frac{\bar{\boldsymbol{\mu}}_{1}\boldsymbol{1}^{T}}{n_{1}}+\frac{\boldsymbol{1}\bar{\boldsymbol{\nu}}_{1}^{T}}{m_{1}}\right),$
$\tilde{\boldsymbol{W}}_{2}$ $=\bar{\lambda}\boldsymbol{1}\boldsymbol{1}^{T}+\gamma\bigg(\frac{\bar{r}_{11}}{m_{1}n_{1}}\boldsymbol{1}\boldsymbol{1}^{T}-\bar{\boldsymbol{A}}(u_{1},v_{1})\bigg).$
4.3.1 A Bound on $\boldsymbol{\tilde{W}_{1}}$
We want to bound the maximum eigenvalue of $\boldsymbol{\tilde{\boldsymbol{W}}_{1}}^{T}\boldsymbol{\tilde{\boldsymbol{W}}_{1}}$ to bound the spectral norm of $\boldsymbol{\tilde{W}_{1}}$. Taking the product $\boldsymbol{\tilde{\boldsymbol{W}}_{1}}^{T}\boldsymbol{\tilde{\boldsymbol{W}}_{1}}$ we have
$\boldsymbol{\tilde{\boldsymbol{W}}_{1}}^{T}\boldsymbol{\tilde{\boldsymbol{W}}_{1}}$ $=\frac{\|\bar{\boldsymbol{\mu_{1}}}\|^{2}}{n_{1}^{2}}\boldsymbol{1}\boldsymbol{1}^{T}+\frac{\boldsymbol{\mu_{1}}^{T}\boldsymbol{1}}{m_{1}n_{1}}\left(\bar{\boldsymbol{\nu_{1}}}\boldsymbol{1}^{T}+\boldsymbol{1}\bar{\boldsymbol{\nu_{1}}}^{T}\right)+\frac{1}{m_{1}}\bar{\boldsymbol{\nu_{1}}}\bar{\boldsymbol{\nu_{1}}}^{T}.$
We will bound each term separately. First, note that
$\bar{\boldsymbol{\mu}}_{1}^{T}\bar{\boldsymbol{\mu}}_{1}=\sum_{i\in U_{1}}\mu_{i}^{2}\leq\|\bar{\boldsymbol{\mu}}_{1}\|_{1}\|\bar{\boldsymbol{\mu}}_{1}\|_{\infty}=\bar{r}_{11}(1-\tilde{\delta})n_{1},$
which implies that the first summand has maximum eigenvalue at most
[…]
since $\|\boldsymbol{1}\boldsymbol{1}^{T}\|=n_{1}.$ By a similar argument, we have
[…]
It remains to bound the remaining term in the decomposition.
Note that $\bar{\boldsymbol{\nu_{1}}}\boldsymbol{1}^{T}+\boldsymbol{1}\bar{\boldsymbol{\nu_{1}}}^{T}$ has rank at most 2, with nontrivial eigenvalues
$\bar{\boldsymbol{\nu}}_{1}^{T}\boldsymbol{1}\pm\|\bar{\boldsymbol{\nu}}_{1}\|_{2}\|\boldsymbol{1}\|_{2}=\bar{\boldsymbol{\nu}}_{1}\boldsymbol{1}^{T}\pm\|\bar{\boldsymbol{\nu}}_{1}\|_{2}\sqrt{n_{1}}.$
We know that
$\bar{\boldsymbol{\nu}}_{1}^{T}\boldsymbol{1}$ $\leq(1-\tilde{\delta})n_{1}$
$\|\bar{\boldsymbol{\nu}}_{1}\|_{2}^{2}$ $\leq\bar{r}_{11}(1-\tilde{\delta})n_{1}.$
Therefore,
$\|\bar{\boldsymbol{\nu}}_{1}\boldsymbol{1}^{T}+\boldsymbol{1}\bar{\boldsymbol{\nu}}_{1}^{T}\|_{2}\leq(1-\tilde{\delta})n_{1}+\sqrt{\bar{r}_{11}(1-\tilde{\delta})n_{1}}\leq 2\bar{r}_{11},$
since $(1-\tilde{\delta})n_{1}\leq\bar{r}_{11}$. This implies that
$\left\|\frac{\bar{\boldsymbol{\mu}}_{1}^{T}\boldsymbol{1}}{m_{1}n_{1}}\bigg(\boldsymbol{1}\bar{\boldsymbol{\nu}}_{1}^{T}+\bar{\boldsymbol{\nu}}_{1}\boldsymbol{1}^{T}\bigg)\right\|\leq\frac{(1-\tilde{\delta})m_{1}}{m_{1}}\|\bar{\boldsymbol{\nu}}_{1}\boldsymbol{1}^{T}+\boldsymbol{1}\bar{\boldsymbol{\nu}}_{1}^{T}\|_{2}\leq 4(1-\tilde{\delta})\bar{r}_{11}.$ (60)
Combining (58), (59), and (60) and applying the triangle inequality shows that
$\|\boldsymbol{\tilde{W}_{1}}\|\leq 2\sqrt{(1-\tilde{\delta})\bar{r}_{11}}.$ (61)
4.3.2 A Bound on $\boldsymbol{\tilde{W}_{2}}$
By construction,
$\boldsymbol{\tilde{W}_{2}}]_{ij}=\begin{cases}2(1-\tilde{\delta})+\frac{\bar{r}_{11}}{m_{1}n_{1}},&\text{if }(i,j)\notin\Omega,\\ 2(1-\tilde{\delta})+\frac{r_{11}}{m_{1}n_{1}},&\text{if }(i,j)\in\Omega.\end{cases}$
Here, we have chosen
$\lambda=\frac{1}{m_{1}n_{1}}+2(1-\tilde{\delta})\gamma.$
Recall that the set $\Omega\cap U_{1}\times V_{1}$ has cardinality $\bar{r}_{11}=m_{1}n_{1}-r_{11}.$ It follows that
$\|\boldsymbol{\tilde{W}_{2}}\|^{2}\leq\|\boldsymbol{\tilde{W}_{2}}\|_{F}^{2}$ $=r_{11}\left(2(1-\tilde{\delta})+\frac{\bar{r}_{11}}{m_{1}n_{1}}\right)+(m_{1}n_{1}-r_{11})\left(2(1-\tilde{\delta})+\frac{r_{11}}{m_{1}n_{1}}\right)$
$=4(1-\tilde{\delta})^{2}m_{1}n_{1}+\frac{r_{11}(m_{1}n_{1}-r_{11})}{m_{1}n_{1}}$
$\leq 4(1-\tilde{\delta})^{2}m_{1}n_{1}+\bar{r}_{11}.$ (62)
4.3.3 Finishing the Bound on $\boldsymbol{W}$
We can obtain the desired bound on $\|\boldsymbol{W}\|$ by applying the triangle inequality and combining the partial bounds given by (55), (56), (61), and (62):
$\|\boldsymbol{W}\|$ $\leq\|\boldsymbol{W_{1}}\|+\|\boldsymbol{W_{2}}\|+\|\boldsymbol{W_{3}}\|$
$\leq\max\left\{\max_{q\geq 2}\lambda\sqrt{r_{qq}},\gamma\left(2\sqrt{(1-\tilde{\delta})\bar{r}_{11}}+\sqrt{4(1-\tilde{\delta})^{2}m_{1}n_{1}+\bar{r}_{11}}\right)\right\}$
$\hskip 28.90755pt+\frac{\lambda}{\sqrt{1-\delta}}\sqrt{r_{1}+r_{2}}+\lambda\sqrt{r_{3}}$ (63)
Note that
$\|\boldsymbol{W_{2}}\|+\|\boldsymbol{W_{3}}\|=\left(\frac{1}{m_{1}n_{1}}+\frac{\tilde{c}(1-\tilde{\delta})}{\sqrt{m_{1}n_{1}}}\right)\left(\sqrt{\frac{r_{1}+r_{2}}{1-\delta}}+\sqrt{r_{3}}\right)<\frac{1}{2}$
if $\max\{r_{1},r_{2},r_{3}\}<cm_{1}n_{1}$ for sufficiently small constant $c$. Similarly, note that
$\lambda\sqrt{r_{qq}}<\frac{1}{2}$
if $r_{qq}<cm_{1}n_{1}$ for sufficiently small $c$ for all $q=2,3,\dots,K$.
Finally,
$\|\boldsymbol{\tilde{W}_{1}}\|+\|\boldsymbol{\tilde{W}_{2}}\|$ $\leq\gamma\left(1+2\sqrt{1-\tilde{\delta}}\right)\sqrt{\bar{r}_{11}}+2(1-\tilde{\delta})\sqrt{m_{1}n_{1}}$
$=\frac{1+2\sqrt{1-\tilde{\delta}}}{2\tilde{\delta}-\delta-1}\sqrt{\frac{\bar{r}_{11}}{m_{1}n_{1}}}+\frac{2(1-\tilde{\delta})}{2\tilde{\delta}-\delta-1}<\frac{1}{2}$
if $\bar{r}_{11}<cm_{1}n_{1}$ for sufficiently small $c$ and $\tilde{\delta}$ sufficiently large. Putting everything together shows that $\|\boldsymbol{W}\|<1$ if (6) holds with sufficiently small $c$.
5 Numerical Experiments
5.1 An Algorithm for Solving (3)
The Alternating Direction Method of Multipliers (ADMM) is a powerful optimization technique particularly well-suited for large-scale and constrained problems. In this section, we present an overview of how ADMM is applied to solve our relaxation of the densest submatrix (3), highlighting its decomposition capabilities and iterative structure that enable efficient computation. Comprehensive details on ADMM can be found in Boyd et al. (2011). The methodology is first derived analytically, followed by a rigorous empirical evaluation using both synthetically generated matrices and real-world datasets representing collaboration and communication networks.
The ADMM algorithm is used to solve optimization problems of the form:
$\begin{array}[]{cl}\min&f(\boldsymbol{x})+g(\boldsymbol{y})\\ \mathrm{s.t.}&\boldsymbol{A}\boldsymbol{x}+\boldsymbol{B}y=\boldsymbol{C}.\end{array}$ (64)
This can be generalized to multi-block problems of the form
$\begin{array}[]{cl}\min&\sum_{i=1}^{k}f(\boldsymbol{x}_{i})\\ \mathrm{s.t.}&\sum_{i=1}^{k}\boldsymbol{A}_{i}\boldsymbol{x}_{i}=\boldsymbol{C}\end{array}$ (65)
That is, the ADMM is used to solve optimization problems with separable objective function subject to linear coupling constraints.
We apply a multi-block ADMM to solve (3) as proposed in Bombina and Ames (2020). We can rewrite the convex program given in (3) to fit the framework (65) by introducing auxiliary variables $\boldsymbol{Q},\boldsymbol{W},\boldsymbol{Z}$:
$\begin{array}[]{cl}\min&\lVert\boldsymbol{X}\rVert_{*}+\gamma\lVert\boldsymbol{Y}\rVert_{1}+\mathbb{I}_{\Omega_{Q}}(\boldsymbol{Q})+\mathbb{I}_{\Omega_{W}}(\boldsymbol{W})+\mathbb{I}_{\Omega_{Z}}(\boldsymbol{Z})\\ \mathrm{s.t.}&\boldsymbol{X}+\boldsymbol{Y}=\boldsymbol{Q},~\boldsymbol{X}=\boldsymbol{W},~\boldsymbol{X}=\boldsymbol{Z},\end{array}$ (66)
where $\Omega_{Q}$, $\Omega_{W}$, and $\Omega_{Z}$ denote the sets
$\Omega_{Q}$ $:=\{\boldsymbol{Q}\in\mathbb{R}^{M\times N}:P_{\tilde{N}}(\boldsymbol{Q})=\boldsymbol{0}\},$
$\quad\Omega_{W}$ $:=\{\boldsymbol{W}\in\mathbb{R}^{M\times N}:\boldsymbol{1}^{T}\boldsymbol{W}\boldsymbol{1}=m_{1}n_{1}\},$
$\Omega_{Z}$ $:=\{\boldsymbol{Z}\in\mathbb{R}^{M\times N}:0\leq Z_{ij}\leq 1\,\forall(i,j)\in\boldsymbol{v}\times\boldsymbol{v}\}.$
Here, $\mathbb{I}_{S}:\mathbb{R}^{M\times N}\to\{0,+\infty\}$ represents the indicator function of the set $\boldsymbol{S}\subseteq\mathbb{R}^{M\times N}$, where $\mathbb{I}_{S}(\boldsymbol{X})=0$ if $\boldsymbol{X}\in\boldsymbol{S}$, and $\infty$ otherwise.
To solve (66), we employ an iterative approach based on the ADMM. In each iteration, we sequentially update each primal decision variable by minimizing the augmented Lagrangian to approximate the dual functional gradient and then apply an approximate dual gradient ascent step.
We define the augmented Lagrangian $\mathcal{L}_{\tau}$ as follows:
$\mathcal{L}_{\tau}$ $=\|\boldsymbol{X}\|_{*}+\gamma\|\boldsymbol{Y}\|_{1}+\mathbb{I}_{\Omega_{Q}}(\boldsymbol{Q})+\mathbb{I}_{\Omega_{W}}(\boldsymbol{W})+\mathbb{I}_{\Omega_{Z}}(\boldsymbol{Z})$
$\quad+\mathrm{Tr}(\boldsymbol{\Lambda}_{Q}^{\top}(\boldsymbol{X}+\boldsymbol{Y}-\boldsymbol{Q}))+\mathrm{Tr}(\boldsymbol{\Lambda}_{W}^{\top}(\boldsymbol{X}-\boldsymbol{W}))+\mathrm{Tr}(\boldsymbol{\Lambda}_{Z}^{\top}(\boldsymbol{X}-\boldsymbol{Z}))$
$\quad+\frac{\tau}{2}\left(\|\boldsymbol{X}+\boldsymbol{Y}-\boldsymbol{Q}\|_{F}^{2}+\|\boldsymbol{X}-\boldsymbol{W}\|_{F}^{2}+\|\boldsymbol{X}-\boldsymbol{Z}\|_{F}^{2}\right),$
where $\tau$ is a regularization parameter selected to ensure that $\mathcal{L}_{\tau}$ is strongly convex in each of the primal variables.
In each iteration, the primal variables are updated sequentially using a Gauss–Seidel strategy to minimize $\mathcal{L}_{\tau}$ with respect to each of $\boldsymbol{X}$, $\boldsymbol{Y}$, $\boldsymbol{Q}$, $\boldsymbol{W}$, and $\boldsymbol{Z}$, followed by updates to the dual variables $\boldsymbol{\Lambda}_{Q}$, $\boldsymbol{\Lambda}_{W}$, and $\boldsymbol{\Lambda}_{Z}$. The updates of the auxiliary variables $\boldsymbol{Q}$, $\boldsymbol{W}$, and $\boldsymbol{Z}$ reduce to orthogonal projections onto the constraint sets $\Omega_{Q}$, $\Omega_{W}$, and $\Omega_{Z}$, respectively; each of these projections can be computed in closed form.
The subproblems for updating $\boldsymbol{X}$ and $\boldsymbol{Y}$ admit closed-form solutions via the elementwise soft-thresholding operator $S_{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined as
$(S_{\phi}(\boldsymbol{x}))_{i}=\begin{cases}x_{i}-\phi,&\text{if }x_{i}>\phi,\\ 0,&\text{if }|x_{i}|\leq\phi,\\ x_{i}+\phi,&\text{if }x_{i}<-\phi.\end{cases}$
It has been established (see Hong and Luo (2017)) that ADMM exhibits linear convergence when applied to the minimization of convex, separable functions under mild regularity assumptions. Accordingly, Algorithm 1 is expected to converge to a global optimum of problem (3). The algorithm is terminated once the primal and dual residuals, $\|\boldsymbol{X}^{(t+1)}-\boldsymbol{W}^{(t+1)}\|_{F},\|\boldsymbol{X}^{(t+1)}-\boldsymbol{Z}^{(t+1)}\|_{F},\|\boldsymbol{W}^{(t+1)}-\boldsymbol{W}^{(t)}\|_{F},\|\boldsymbol{Q}^{(t+1)}-\boldsymbol{Q}^{(t)}\|_{F},\|\boldsymbol{Z}^{(t+1)}-\boldsymbol{Z}^{(t)}\|_{F},\|\boldsymbol{\Lambda}_{Q}^{(t+1)}-\boldsymbol{\Lambda}_{Q}^{(t)}\|_{F},\|\boldsymbol{\Lambda}_{W}^{(t+1)}-\boldsymbol{\Lambda}_{Z}^{(t)}\|_{F}$ and $\|\boldsymbol{\Lambda}_{Z}^{(t+1)}-\boldsymbol{\Lambda}_{Q}^{(t)}\|_{F}$, are all within desired error bounds. The algorithmic steps are summarized in Algorithm 1.
Input: Adjacency matrix $\boldsymbol{A}\in\mathbb{R}^{M\times N}$, Target submatrix size $(m,n)$, Regularization parameter $\gamma$, Augmented Lagrangian penalty $\tau$, Stopping tolerance $\epsilon$, Maximum iterations maxiter.
Output: Primal variables $\boldsymbol{X},\boldsymbol{Y},\boldsymbol{Q}$, number of iterations iter.
Initialization:
Set $\mu=1/\tau$ (reciprocal penalty parameter).
Initialize $\boldsymbol{W}=\boldsymbol{X}=\boldsymbol{Y}=\boldsymbol{Z}=\texttt{constant matrix}$ scaled to match $mn/MN$.
Set dual variables $\boldsymbol{\Lambda}_{Q}=\boldsymbol{\Lambda}_{Z}=\boldsymbol{\Lambda}_{W}=\boldsymbol{0}$.
Set $\texttt{iter}=0$; convergence $\texttt{false}$.
while convergence == false
   Increment iteration counter: $\texttt{iter}\leftarrow\texttt{iter}+1$.
 Update $\boldsymbol{Q}$:
 $\boldsymbol{Q}\leftarrow(\boldsymbol{X}-\boldsymbol{Y}+\mu\boldsymbol{\Lambda}_{Q})$
   Elementwise mask: $\boldsymbol{Q}\leftarrow\boldsymbol{Q}\circ\boldsymbol{A}$ (Hadamard product)
 Update $\boldsymbol{X}$:
 $\boldsymbol{X}\leftarrow\texttt{MatrixShrink}\left(\frac{1}{3}(\boldsymbol{Y}+\boldsymbol{Q}+\boldsymbol{Z}+\boldsymbol{W}-\mu(\boldsymbol{\Lambda}_{Q}+\boldsymbol{\Lambda}_{W}+\boldsymbol{\Lambda}_{Z})),\frac{1}{3\tau}\right)$
 Update $\boldsymbol{Y}$:
 $\boldsymbol{Y}\leftarrow\max\left(\boldsymbol{X}-\boldsymbol{Q}-\gamma\mu\boldsymbol{1}\boldsymbol{1}^{T}+\mu\boldsymbol{\Lambda}_{Q},\boldsymbol{0}\right)$
 Update $\boldsymbol{W}$:
 $\boldsymbol{W}\leftarrow\boldsymbol{X}+\mu\boldsymbol{\Lambda}_{W}$
   Shift to enforce $\sum W_{ij}=mn$ by adding constant offset.
 Update $\boldsymbol{Z}$:
 $\boldsymbol{Z}\leftarrow\texttt{clip}(\boldsymbol{X}+\mu\boldsymbol{\Lambda}_{Z},\boldsymbol{0},\boldsymbol{1}\boldsymbol{1}^{T})$(clip between 0 and 1)
 Update Dual Variables:
 $\boldsymbol{\Lambda}_{Q}\leftarrow\boldsymbol{\Lambda}_{Q}+\tau(\boldsymbol{X}-\boldsymbol{Y}-\boldsymbol{Q})$
 $\boldsymbol{\Lambda}_{W}\leftarrow\boldsymbol{\Lambda}_{W}+\tau(\boldsymbol{X}-\boldsymbol{W})$
 $\boldsymbol{\Lambda}_{Z}\leftarrow\boldsymbol{\Lambda}_{Z}+\tau(\boldsymbol{X}-\boldsymbol{Z})$
 Check Convergence:
   Primal residual: $\epsilon_{P}=\max\left\{\|\boldsymbol{X}-\boldsymbol{Z}\|_{F},\|\boldsymbol{X}-\boldsymbol{W}\|_{F},\|\boldsymbol{X}-\boldsymbol{Y}-\boldsymbol{Q}\|_{F}\right\}/\|\boldsymbol{X}\|_{F}$
   Dual residual: $\epsilon_{D}=\max\left\{\|\boldsymbol{Z}-\boldsymbol{Z}_{\text{old}}\|_{F},\|\boldsymbol{W}-\boldsymbol{W}_{\text{old}}\|_{F},\|\boldsymbol{Q}-\boldsymbol{Q}_{\text{old}}\|_{F}\right\}/\|\boldsymbol{X}\|_{F}$
 if $\epsilon_{P}<\epsilon$ and $\epsilon_{D}<\epsilon$ or $\texttt{iter}\geq\texttt{maxiter}$
      Set convergence = true.
   end if
end while
return $\boldsymbol{X},\boldsymbol{Y},\boldsymbol{Q},\texttt{iter}$.
Algorithm 1 DENSUB: ADMM for Densest Submatrix/Subgraph Recovery
5.2 Analysis of Synthetic Data
We perform two experiments using matrices sampled from the planted submatrix model to illustrate the theoretical sufficient conditions for perfect recovery of the densest submatrix given by Theorem 2.1. In each, we consider random matrices sampled from the planted submatrix model while varying problem parameters such as the density and size of the planted blocks, and record the number of times that we correctly identify a particular submatrix from the solution of (3). This allows us to empirically model the probability of recovery of the hidden submatrix for each problem setting and, hence, illustrate the phase transitions to perfect recovery given in Theorem 2.1 and Theorem 2.2.
We have prepared Python [^1] https://github.com/bpames/DENSUB, MATLAB [^2] https://github.com/pbombina/admmDSM, and R [^3] https://cran.r-project.org/web/packages/admmDensestSubmatrix/index.html implementations of Alg. 1. We use the Python implementation in the following experiments. All experiments were completed using a single node of the University of Southampton’s Iridis 5 HPC cluster.
5.2.1 Experiment 1
We first consider $500\times 500$ matrices sampled from the planted dense submatrix model with $k=\lfloor M/m\rfloor$ blocks constructed as follows for each $(q,m)$ pairs with
$q$ $\in\{0.30,0.35,0.4,0.45,0.5,0.55,0.6,0.65,0.7,0.75,0.85,0.9,0.95,1\},$
$m$ $\in\{50,75,100,125,150,175,200,225,250\}:$
  • We partition the rows into k1k-1 sets $U_{1},U_{2},\dots,U_{k-1}$ of size $m$. The remaining M(k1)mM-(k-1)m rows are assigned to remaining block $U_{k}$. We partition the columns symmetrically.
  • We sample $\boldsymbol{A}$ from the symmetric planted dense submatrix model according to this partition with probabilities
    $p_{11}$ $=q,$
    $p_{jj}$ $=\frac{1}{4(j-1)}\;\;\text{for }j=2,3,\dots,k,$
    $p_{ij}$ $=p_{ii}p_{jj}\;\;\text{if }i\neq j.$
    Here, we sample $\boldsymbol{A}$ symmetrically so that $a_{ij}=a_{ji}$ for all $i,j\in[M]=[500]$.
This construction ensures that the $(U_{1},U_{1})$ block is the densest $(m,m)$-submatrix in expectation. Moreover, Conditions (4a) and (4b) hold by our choice of $m$ and $M$. If the gap $q-p_{11}$ is sufficiently large then (4c) is satisfied and we should expect that the matrix representation of $(U_{1},U_{1})$ is the unique optimal solution for (3). When this gap is small, i.e., $q$ is close to $1/2$, then we may not have recovery w.h.p. because the sufficient conditions given by Theorem 2.1 are not satisfied.
This is illustrated by the examples given in Figure 2. Each of the matrices $\boldsymbol{A}$ have a densest block indexed by $(U_{1},U_{1})=([m],[m])$ for $m=100$. The remaining four diagonal blocks are less dense and the off-diagonal blocks are relatively sparse. We can create an undirected graph $G$ with adjacency matrix equal to each sampled matrix $\boldsymbol{A}$. These graphs have densest $m$-subgraphs indexed by $U_{1}$ in expectation. When the gap between $q$ and $p_{22}$, the density of the next densest diagonal block, is small, we have two planted $(m,m)$-submatrices with very similar densities. We cannot expect to recover the densest $m$-subgraph or densest $(m,m)$-submatrix in this case (see Figs. 2(c) and 2(d)). However, when the gap is larger, the corresponding graph $G$ consists of a single dense $m$-subgraph obscured by sparse noise in the form of diversionary edges and nodes. In this case, we should expect to identify the planted dense submatrix and subgraph from the solution of (3) (see Figs. 2(a) and Figs. 2(b)).
To empirically verify this predicted behaviour, we generate $10$ problem instances $\boldsymbol{A}$ corresponding to each $(q,m)$-pair. For each of these matrices, we solve (3) with
$\gamma=\frac{6}{m(q-\min\{p_{ij}\})}$
using the Python implementation of Alg. 1. We use the augmented Lagrangian parameter $\mu=2$, stopping tolerance $10^{-4}$, and maximum number of iterations $2000$ in each call to the solver. We round the entries of $\boldsymbol{X}$ to the nearest integer and compare to the matrix representation of the planted submatrix indexed by $(U_{1},U_{1})$. We say that we have recovered the planted submatrix if
$\frac{\|\boldsymbol{X}-\boldsymbol{X}_{0}\|_{F}}{\|\boldsymbol{X}_{0}\|_{F}}<10^{-3}.$
We should expect a sharp transition from no recoveries in the 10 trials to recovery in all 10 trials as both $m$ and $q$ increase, as suggested by the sufficient condition (4c).
The results of this experiment are summarized in Figure 4(a). The observed recovery rates closely match that predicted by Theorem 2.1. Indeed, when $q$ and $m$ satisfy (4c), the solution computed by Alg. 1 is exactly the matrix representation of the planted submatrix in almost all trials.
(a) Easy Problem $\boldsymbol{A}$
(b) Easy Problem $G$
(c) Hard Problem $\boldsymbol{A}$
(d) Hard Problem $G$
Figure 2: Examples of easy and hard to solve instances of the densest submatrix problem sampled in Experiment 1. Here, orange squares indicate the nodes $U_{1}$ in the graph $G$ with adjacency matrix $\boldsymbol{A}$ that induce the planted subgraph.
5.2.2 Experiment 2
We next consider problem instances of the densest $(m,m)$-submatrix problem constructed as follows for each $(q,m)$-pair in
$q$ $\in\{0.25,0.30,0.35,0.4,0.45,0.5,0.55,0.6,0.65,0.7,0.75,0.85,0.9,0.95,1\}$
$m$ $\in\{200,225,250,275,300,325,350,375\}.$
  • We divide $[M]=[500]$ into $K=2$ blocks $U_{1},U_{2}$ of size $m$ and Mm=500mM-m=500-m, respectively.
  • We sample matrix $\boldsymbol{A}$ from the planted dense submatrix model with probabilities
    $p_{11}=p_{22}$ $=q$
    $p_{12}=p_{21}$ $=0.25.$
By construction, $(U_{1},U_{1})$ indexes the densest $(m,m)$-submatrix and $U_{1}$-subgraph of the corresponding undirected graph $G$ when $m>M/2$. If $m<M/2$ then the matrix contains several $(m,m)$-submatrices as dense as that indexed by $(U_{1},U_{1})$; in this case, it is impossible to guarantee that we will recover the matrix representation of $(U_{1},U_{1})$ as the optimal solution of (3); see Figs. 3(g) and 3(h) for an example of such a matrix and graph. On the other hand, the sufficient conditions for recovery given by Theorem 2.1 and (4c), in particular, may not hold if
  • $q\approx 0.25$, as illustrated by Figs. 3(c) and 3(d); or
  • $m\approx M-m$, as illustrated by Figs. 3(e) and 3(f).
In the first case, the density in the diagonal blocks is close to that of the off-diagonal blocks and we should not expect to distinguish between them. In the second, the sizes of the two diagonal blocks are close and, again, we should not expect to recover $(U_{1},U_{1})$ as the index set of the densest $(m,m)$-submatrix. However, the sufficient conditions given by Theorem 2.1 suggest that we will if have perfect recovery of $(U_{1},U_{1})$ from the optimal solution of (3) w.h.p. if $m$ and $q$ are sufficiently large (see Figs. 3(a) and 3(b)).
As before, we empirically verify the phase transition to perfect recovery by generating $10$ problem instances $\boldsymbol{A}$ corresponding to each $(q,m)$-pair. For each of these matrices, we solve (3) with
$\gamma=\frac{6}{m(q-p_{12})}$
using the Python implementation of Alg. 1. We use the augmented Lagrangian parameter $\tau=2$, stopping tolerance $10^{-4}$, and maximum number of iterations $2000$ as in the previous experiment. We record each computed solution $\boldsymbol{X}$ after rounding and compare to the matrix representation of the planted submatrix indexed by $(U_{1},U_{1})$. We say that we have recovered the planted submatrix if
$\frac{\|\boldsymbol{X}-\boldsymbol{X}_{0}\|_{F}}{\|\boldsymbol{X}_{0}\|_{F}}<10^{-3}$
and count the number of recoveries of $\boldsymbol{X}_{0}$ for each $(q,m)$-pair.
The results of this experiment are summarized in Figure 4(b). As in Section 5.2.1, we have empirical recovery rates that match those predicted by Theorem 2.1. Again, when $q$ and $m$ satisfy (4c) we have perfect recovery of $\boldsymbol{X}_{0}$ in almost all trials. The reported recovery rates are perhaps overly pessimistic in this case. We record recovery of the submatrix indexed by $(U_{1},U_{1})$. When Mm>>M-m>>, this submatrix is not the unique densest $(m,m)$-submatrix in expectation. In this case, the solution $\boldsymbol{X}$ is a convex combination of several $(m,m)$-submatrix of similar densities. We will see in the next section that we can identify the submatrices comprising these solutions using simple rounding techniques in certain cases.
(a) Easy Problem $\boldsymbol{A}$
(b) Easy Problem $G$
(c) Hard Problem $\boldsymbol{A}$
(d) Hard Problem $G$
(e) Another Hard $\boldsymbol{A}$
(f) Another Hard $G$
(g) Impossible Recovery $\boldsymbol{A}$
(h) Impossible Recovery $G$
Figure 3: Examples of easy, hard, and impossible graphs for recovery of the planted submatrix sampled in Experiment 2. Here, orange squares indicate the nodes $U_{1}$ in the graph $G$ with adjacency matrix $\boldsymbol{A}$ that induce the planted subgraph.
(a) Recovery Counts for Experiment 1
(b) Recovery Counts for Experiment 2
Figure 4: Recovery counts for Experiments 1 and 2. Darker squares indicate more recoveries out of 10 trials for each $(q,m)$-pair.
5.3 Identifying Maximum Cliques in Real-World Networks
5.3.1 Benchmarking Networks
We apply Alg. 1 to analyse several widely studied benchmarking network data sets. The maximum clique has been previously identified for each of these networks and is known. We will illustrate that Alg. 1 recovers the maximum clique for each of the following networks:
  • The Jazz Collaboration Network Gleiser and Danon (2003) is a collaboration network where two musicians are connected if they have performed together. It contains a single maximum clique of size $30$.
  • Zachary’s Karate Club Zachary (1977) is a social network of 34 members of a karate club in the 1970s. It contains two maximum cliques of size $5$.
  • Dolphins Lusseau et al. (2003) is a social network of 62 bottlenose dolphins living of Doubtful Sound, New Zealand. It contains three maximum cliques of size $5$.
  • Les Miserables Knuth (1993) is a co-appearance network of characters in the novel Les Miserables containing two maximum cliques of length $10$.
For each network, we apply our Python implementation of Alg. 1 to solve the corresponding instance of (3) with $\gamma=12/m$, where $m=n$ is chosen to be the known size of the maximum clique in the network. We use augmented Lagrangian parameter $\tau=2$ and stop the iterative process if an $\epsilon=10^{-4}$ suboptimal solution is found or $2000$ iterations have been performed.
For each of these networks, we successfully identify the maximum clique(s) present within the network. Indeed, the solution $\boldsymbol{X}$ returned by Alg. 1 when applied to the Jazz Collaboration Network is exactly the rank-one matrix representation of the maximum clique in this graph. On the other hand, each of the remaining networks contain multiple maximum cliques. In each case, the solution $\boldsymbol{X}$ returned by Alg. 1 is a convex combination of the matrix representations of each maximum clique. A simple rounding process, e.g., taking the $m$ largest nonzero entries of the diagonal of $\boldsymbol{X}$ reveals one of the maximum cliques for each of these networks; we can easily identify the remaining cliques by inspection using the clique found in this rounding step. Figure 5 provides a visualisation of these results. For each of these networks, we give a visualisation of the adjacency matrix, the optimal solution of (3) found by Alg. 1, and the maximum clique found by the rounding procedure described above.
(a) Adjacency matrix of JAZZ
(b) Recovered Solution $\boldsymbol{X}$ for JAZZ
(c) Recovered clique in JAZZ
(d) Adjacency matrix of KARATE
(e) Solution $\boldsymbol{X}$ for KARATE
(f) Recovered clique in KARATE
(g) Adjacency matrix of DOLPHINS
(h) Solution $\boldsymbol{X}$ for DOLPHINS
(i) Recovered clique in DOLPHINS
(j) Adjacency matrix of LESMIS
(k) Solution $\boldsymbol{X}$ for LESMIS
(l) Recovered clique in LESMIS
Figure 5: Adjacency matrices of the JAZZ, KARATE, DOLPHINS, and LESMIS networks and recovered maximum cliques.
6 Maximum Cliques of the A Song of Ice and Fire Series
We conclude by applying Alg. 1 to identify the largest clique in each of the 5 novels in George R.R. Martin’s A Song of Ice and Fire (ASOIAF) saga using the character interaction networks considered in Beveridge and Shan (2016); Each network, one for each book in series, has nodes corresponding to characters and edge weights corresponding to the number of times two character names appear within 15 words of one another, which models the number of interactions between each pair of characters in the book. A detailed discussion of the properties of these interaction networks, as well as all network data are available at 37. A similar analysis of the television adaptation of ASOIAF, HBO’s Game of Thrones, is presented in Beveridge and Chemers (2018).
We perform the following preprocessing steps to the character interaction network for each book:
  • We first binarize the network by setting any edge weight greater than a pre-chosen threshold $t$ equal to $1$ and all other edge weights equal to 0. This yields an undirected, unweighted character interaction network, where two characters are adjacent if they have at least $t$ interactions in a given book. We use threshold $t=5$ for Book 5, and threshold $t=10$ for remaining the books.
  • We then form a final interaction network $G$ where two characters are adjacent if there is at least one walk of length $2$ between them in the binarized interaction network.
For each interaction network $G$, we use the find_cliques [^4] https://networkx.org/documentation/stable/reference/algorithms/generated/networkx.algorithms.clique.find_cliques.html function from the NetworkX library to calculate the set of all maximal clique of $G$ using the algorithms developed in Bron and Kerbosch (1973); Tomita et al. (2006); Cazals and Karande (2008); after calculating all maximal cliques, we find the maximum clique as the maximal clique with largest cardinality. We then apply the Python implementation of Alg. 1 to solve the corresponding instance of (3) with the adjacency matrix of $G$ as input and $\gamma=12/m$, where $m=n$ is equal to size of the maximum clique within the character interaction network $G$. As in the experiments in Sect. 5.3, we employ augmented Lagrangian parameter $\tau=2$, stopping tolerance $10^{-4}$, and a maximum of $2000$ iterations.
We correctly identify the unique maximum clique in the interaction network $G$ for Books 1, 2, 3, and 5. Book 4 has two overlapping maximum cliques of size $5$, and we recover a convex combination $\boldsymbol{X}$ of the matrix representations of these cliques; the rounding procedure described in Sect. 5.3 correctly identifies the two maximum cliques in the interaction network for Book 4.
The maximum clique in each interaction network represents the largest collection of fully connected characters in the corresponding book. Here, the membership of these cliques reflects the dynamics of the series of novels. In the first book, we observe a single large community primarily comprised of the Baratheon, Stark, and Lannister families. As the saga progresses, the action of the novels scatters these characters across many disparate locations. This corresponds to interaction networks consisting of much smaller cliques and clusters. For each book and interaction network, we successfully identified the largest of these cliques, but further analysis is required to find all of the underlying communities. This analysis aligns with the more detailed analysis of the network dynamics of the ASOIAF series found in 10; 9; 37.
(a) Adjacency matrix for Book 1
(b) Solution $\boldsymbol{X}$ for Book 1
(c) Recovered clique for Book 1
(d) Adjacency matrix for Book 2
(e) Solution $\boldsymbol{X}$ for Book 2
(f) Recovered clique for Book 2
(g) Adjacency matrix for Book 3
(h) Solution $\boldsymbol{X}$ for Book 3
(i) Recovered clique for Book 3
(j) Adjacency matrix for Book 5
(k) Solution $\boldsymbol{X}$ for Book 5
(l) Recovered clique for Book 5
Figure 6: Adjacency matrices of the character interaction networks for Books 1,2,3, and 5 of the ASOIAF saga and recovered maximum cliques.
(a) Adjacency matrix for Book 4
(b) Solution $\boldsymbol{X}$ for Book 4
(c) Maximum clique for Book 4
(d) Another Maximum clique for Book 4
Figure 7: Adjacency matrix of the character interaction network for Book 4 of the ASOIAF saga with optimal solution $\boldsymbol{X}$ of (3). We obtain two maximum cliques in this network from $\boldsymbol{X}$.
Book $\omega(G)$ Maximum Clique(s)
1 18 Jon-Arryn, Varys, Barristan-Selmy, Pycelle, Stannis-Baratheon, Jaime-Lannister,
Tywin-Lannister, Petyr-Baelish, Renly-Baratheon, Sansa-Stark, Robert-Baratheon,
Eddard-Stark, Bran-Stark, Jon-Snow, Tyrion-Lannister, Catelyn-Stark,
Joffrey-Baratheon, Cersei-Lannister
2 11 Stannis-Baratheon, Jaime-Lannister, Joffrey-Baratheon, Sansa-Stark,
Robb-Stark, Renly-Baratheon, Tyrion-Lannister, Cersei-Lannister, Varys,
Sandor-Clegane, Petyr-Baelish
3 8 Kevan-Lannister, Jaime-Lannister, Sansa-Stark, Joffrey-Baratheon,
Tywin-Lannister, Robb-Stark, Cersei-Lannister, Tyrion-Lannister
4 5 Tommen-Baratheon, Tyrion-Lannister, Tywin-Lannister, Robert-Baratheon,
Osmund-Kettleblack
Tommen-Baratheon, Loras-Tyrell, Osney-Kettleblack, High-Sparrow, Taena-of-Myr
5 11 Daario-Naharis, Skahaz-mo-Kandaq, Missandei, Reznak-mo-Reznak, Belwas,
Hizdahr-zo-Loraq, Barristan-Selmy, Grey-Worm, Quentyn-Martell, Galazza-Galare,
Daenerys-Targaryen
Table 1: Maximum cliques and size of maximum clique $\omega(G)$ in the character interaction network for each book in the ASOIAF series. We report both maximum cliques identified for Book 4.
7 Conclusions
We have extended existing convex relaxation approaches for the densest submatrix and planted clique problems to a significantly more general and realistic setting involving multiple dense substructures. By analyzing both probabilistic and adversarial models, we established new sufficient conditions guaranteeing perfect recovery of a planted submatrix in polynomial time. These conditions reveal the precise signal-to-noise trade-offs that govern the phase transition from partial to exact recovery. Further, our results generalize and strengthen several existing recovery guarantees for planted cliques, bicliques, and dense subgraphs as special cases.
Our results demonstrate that nuclear-norm-based convex formulations can effectively identify planted dense blocks even in heterogeneous stochastic block models and in the presence of structured adversarial noise. The accompanying dual-certificate analysis provides rigorous theoretical justification for these recovery guarantees. Empirical experiments conducted on both synthetic and real-world networks substantiate the theoretical predictions, demonstrating the sharpness of the proposed phase transitions and the robustness of the convex relaxation under heterogeneous noise.
The results of our empirical analysis suggest several areas for future research. Of particular interest is the need to strengthen our sufficient conditions for perfect recovery for the case where we have several planted dense blocks of the same size and similar density, e.g., when graphs contain several maximum cliques. We have observed that our nuclear-norm-based approach often recovers a convex combination of the multiple optimal solutions, but we have yet to identify a theoretical phase transition ensuring perfect recovery in this case.
Moreover, it is necessary to develop scalable algorithmic implementations tailored for large-scale network data. The ADMM approach considered in Section 5 is considerably more efficient than classical interior point methods for solving our convex relaxation, but still rely on a computationally expensive singular value decomposition, which limits the scale of problems that we can solve in practice. Scalable first-order methods that avoid this SVD step or efficient interior-point algorithms exploiting the sparse structure of the constraints are needed to obtain scalable algorithms for this problem.
Acknowledgements
The authors acknowledge the use of the IRIDIS High Performance Computing Facility, and associated support services at the University of Southampton, in the completion of this work. B. Ames was supported in part by National Science Foundation Grants #20212554 and #2108645, as well as University of Alabama Research Grants RG14678 and RG14838.
Appendix A Proof of Lemma 3.5
Lemma 3.5 is a special case of the Matrix Bernstein Inequality applied to $\boldsymbol{Z}={\boldsymbol{\Theta}}-\tilde{\boldsymbol{\Theta}}$ as found in the hypothesis of Lemma 3.5. The Matrix Bernstein Inequality, as presented in [44], is provided below:
Theorem A.1.
Let $\{\boldsymbol{S}_{k}\}$ be a finite sequence of independent real $d_{1}\times d_{2}$ random matrices such that $\textbf{E}[\boldsymbol{S}_{k}]=\boldsymbol{0}$ and $\lVert\boldsymbol{S}_{k}\rVert\leq L$ for all $k$. Introduce the sum $\boldsymbol{Z}:=\sum_{k}\boldsymbol{S}_{k}$ and let $\nu(\boldsymbol{Z})$ denote the variance of the sum:
$\nu(\boldsymbol{Z})=\max\{\lVert\mathbf{E}[\boldsymbol{Z}\boldsymbol{Z}^{T}]\rVert,\lVert\mathbf{E}[\boldsymbol{Z}^{T}\boldsymbol{Z}]\rVert\}$ $=\max\left\{\left\lVert\sum_{k}\mathbf{E}[\boldsymbol{S}_{k}\boldsymbol{S}_{k}^{T}]\right\rVert,\left\lVert\sum_{k}\mathbf{E}[\boldsymbol{S}_{k}^{T}\boldsymbol{S}_{k}]\right\rVert\right\}.$ (67)
Then, for all $t\geq 0$,
$\operatorname{Pr}(\lVert\boldsymbol{Z}\rVert\geq t)$ $\leq(d_{1}+d_{2})\exp\left(\frac{-t^{2}/2}{\nu(\boldsymbol{Z})+Lt/3}\right).$ (68)
Consider the matrix
$\boldsymbol{S}_{j}:=\boldsymbol{d}_{j}\boldsymbol{e}_{j}^{T},$ (69)
where $\boldsymbol{d}_{j}:=\boldsymbol{a}_{j}-\boldsymbol{\tilde{a}}_{j}$ is the difference of the $j$th columns $\boldsymbol{a}_{j}$, $\boldsymbol{\tilde{a}}_{j}$ of $\boldsymbol{\Theta}$ and $\tilde{\boldsymbol{\Theta}}$, respectively, and $\boldsymbol{1}_{j}$ denotes the $j$th column of the identity matrix. Clearly $\boldsymbol{Z}=\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}=\sum_{j}\boldsymbol{S}_{j}$. We next calculate $L$ and $\nu(\boldsymbol{Z})$ to substitute in (68). We start with a bound on $L$.
We start with the following lemma, which bounds $\|\boldsymbol{S}_{j}\|$ with high probability.
Lemma A.1.
There exist $c_{1}\geq 0$ such that matrices $\boldsymbol{S}_{j}$ defined by (69) satisfy
$\|\boldsymbol{S}_{j}\|\leq L:=c_{1}\sqrt{\max\{\tilde{\sigma}_{*}^{2},1\}\log N}$ (70)
for all $j=1,2,\dots N$ with high probability, where
$\tilde{\sigma}_{*}^{2}:=\max_{s=1,2,\dots,K}\tilde{\sigma}_{s}^{2}=\max_{s=1,2,\dots,K}\frac{p_{s}}{1-p_{s}}.$ (71)
Proof.
Fix $j\in V_{s}$ for some $s=1,2,\dots,K$. Then,
$\lVert\boldsymbol{S}_{j}\rVert=\lVert\boldsymbol{d}_{j}\boldsymbol{e}^{T}_{j}\rVert=\lVert\boldsymbol{d}_{j}\rVert\lVert\boldsymbol{e}^{T}_{j}\rVert=\lVert\boldsymbol{d}_{j}\rVert.$
We have
$\lVert\boldsymbol{d}_{j}\lVert^{2}$ $\leq(n-n_{j})\left(-\frac{n_{j}}{n-n_{j}}+\frac{p_{s}}{1-p_{s}}\right)^{2}=\frac{(n_{j}-np_{s})^{2}}{(1-p_{s})^{2}(n-n_{j})}.$
Note further that
$\frac{d}{dn_{j}}\|\boldsymbol{d}_{j}\|^{2}$ $=\frac{(np_{s}-n_{j})(n_{j}-n(2-p_{s}))}{(n-n_{j})^{2}(1-p_{s})^{2}},$
$\frac{d^{2}}{d^{2}n_{j}}\|\boldsymbol{d}_{j}\|^{2}$ $=\frac{2n^{2}}{(n-n_{j})^{3}}.$
The second derivative is strictly positive for all $n_{j}<n$, which implies that $\|\boldsymbol{d}_{j}\|^{2}$ is strictly convex with respect to $n_{j}$ on the interval $[0,n)$. Moreover, $\|\boldsymbol{d}_{j}\|^{2}$ is strictly decreasing on $0,p_{s}n)$ and strictly increasing on $(p_{s}n,n)$. Thus, $\|\boldsymbol{d}_{j}\|^{2}$ is minimized at the stationary point $n_{j}=p_{s}n$ and is maximized at one of the two endpoints $a$ and $b$ when restricted to the interval $[a,b]$ for any $0<a<p_{s}n<b<n$.
Recall that applying the standard Bernstein inequality (33) shows that
$p_{s}n-6\max\left\{\sqrt{\sigma_{s}^{2}n\log N},\log N\right\}<n_{j}<p_{s}n-6\max\left\{\sqrt{\sigma_{s}^{2}n\log N},\log N\right\}$
with high probability. Substituting these endpoints as $n_{j}$ in the formula for $\|\boldsymbol{d}_{j}\|^{2}$ shows that if $n-n_{j}\geq\bar{c}>0$ then
$\lVert\boldsymbol{d}_{j}\rVert^{2}$ $\leq\frac{36\max\{n\sigma_{p_{s}}^{2}\log{N},(\log{N})^{2}\}}{(1-p_{s})^{2}(n-n_{j})}$
$\leq\frac{36\bar{c}\max\{n\sigma_{p_{s}}^{2}\log{N},(\log{N})^{2}\}}{(1-p_{s})^{2}n}$
$=\mathcal{O}\left(\max\{\tilde{\sigma}_{s}^{2},1\}\log N\right)=\mathcal{O}\left(\max\{\tilde{\sigma}_{*}^{2},1\}\log N\right)$
w.h.p., where the final inequality follows from the assumption $\log N<n(1-p_{s})^{2}$, which implies that
$\frac{(\log N)^{2}}{n(1-p_{s})^{2}}\leq\log N.$
The existence of such a $\bar{c}>0$ is guaranteed w.h.p. by the Bernstein inequality. Applying the union bound over all $j$ shows that
$\|\boldsymbol{S}_{j}\|\leq L:=\mathcal{O}\left(\sqrt{\max\{\tilde{\sigma}_{*}^{2},1\}\log N}\right)$
with high probability. ∎∎
The next lemma provides a bound on $\nu(\boldsymbol{Z})$.
Lemma A.2.
The matrix $\boldsymbol{Z}=\sum_{j}\boldsymbol{S}_{j}$ defined by (69) satisfies
$\nu(\boldsymbol{Z})\leq\bar{c}\sum_{s=1}^{K}\tilde{\sigma}_{s}^{2}n_{s}\leq\bar{c}~\tilde{\sigma}_{*}^{2}N$
with high probability for any constant $\bar{c}>0$ satisfying $n-n_{j}\geq\frac{n}{\bar{c}}$ for all $j$.
Proof.
Constructing upper bounds for both $\lVert\mathbf{E}[\boldsymbol{Z}\boldsymbol{Z}^{T}]\rVert$ and $\lVert\mathbf{E}[\boldsymbol{Z}^{T}\boldsymbol{Z}]\rVert$ suffices. Recall that $\boldsymbol{Z}=\sum_{j}\boldsymbol{S}_{j}=\sum_{j}\boldsymbol{d}_{j}\boldsymbol{e}_{j}^{T}$. By our choice of $\boldsymbol{Z}$, we have
$\boldsymbol{Z}^{T}\boldsymbol{Z}=\sum_{i=1}^{N}(\boldsymbol{d}_{i}\boldsymbol{e}_{i}^{T})^{T}\sum_{j=1}^{N}(\boldsymbol{d}_{j}\boldsymbol{e}_{j}^{T})=\sum_{i=1}^{N}\sum_{j=1}^{N}\left(\boldsymbol{d}_{i}^{T}\boldsymbol{d}_{j}\right)\boldsymbol{e}_{i}\boldsymbol{e}_{j}^{T}.$
The independence of $\boldsymbol{d}_{i}$, $\boldsymbol{d}_{j}$ and the fact that $\textbf{E}[\boldsymbol{d}_{j}]=\boldsymbol{0}$ for all $i,j$ implies that that $\mathbf{E}[\boldsymbol{Z}^{T}\boldsymbol{Z}]$ is a diagonal matrix with $j$th diagonal entry equal to $\textbf{E}\|\boldsymbol{d}_{j}\|^{2}$. Consequently,
$\lVert\mathbf{E}[\boldsymbol{Z}^{T}\boldsymbol{Z}]\rVert=\max_{j}\mathbf{E}\lVert\boldsymbol{d}_{j}\rVert^{2}.$
For each $j\in V_{s}$, we have
$\mathbf{E}\lVert\boldsymbol{d}_{j}\rVert^{2}$ $=\mathbf{E}\left[\frac{(n_{j}-np_{s})^{2}}{(1-p_{s})^{2}(n-n_{j})}\right].$
Recall that $n-n_{j}\geq n/\bar{c}$ for some constant $\bar{c}>0$ with high probability by the Bernstein bound. This implies that
$\mathbf{E}\lVert\boldsymbol{d}_{j}\rVert^{2}$ $\leq\bar{c}~\mathbf{E}\left[\frac{(n_{j}-np_{s})^{2}}{(1-p_{s})^{2}n}\right]=\frac{\bar{c}\operatorname{Var}(n_{j})}{(1-p_{s})^{2}n}$
$=\frac{\bar{c}~p_{s}}{1-p_{s}}=\bar{c}~\tilde{\sigma}_{s}$ (72)
with high probability, since $n_{j}$ is binomially distributed with variance $p_{s}(1-p_{s})n$.
On the other hand, the triangle inequality and Jensen’s inequality imply that
$\lVert\mathbf{E}[\boldsymbol{Z}\boldsymbol{Z}^{T}]\rVert$ $\leq\sum_{j\notin V_{1}}\|\mathbf{E}[\boldsymbol{d}_{j}\boldsymbol{d}_{j}^{T}]\|$
$=\sum_{s=1}^{k}\sum_{j\in V_{s}}\|\mathbf{E}[\boldsymbol{d}_{j}\boldsymbol{d}_{j}^{T}]\|$
$\leq\sum_{s=1}^{k}\sum_{j\in V_{s}}\mathbf{E}\|\boldsymbol{d}_{j}\|^{2}$
$\leq\bar{c}\sum_{s=1}^{k}\left(\frac{p_{s}}{1-p_{s}}\right)n_{s}=\bar{c}\sum_{s=1}^{k}\tilde{\sigma}_{s}^{2}n_{s}$ (73)
with high probability. Combining (72) and (73) yields
$\nu(\boldsymbol{Z})=\max\Big\{\lVert\mathbf{E}[\boldsymbol{Z}^{T}\boldsymbol{Z}]\rVert,\lVert\mathbf{E}[\boldsymbol{Z}\boldsymbol{Z}^{T}]\rVert\Big\}=\bar{c}\sum_{s=1}^{k}\tilde{\sigma}_{s}^{2}n_{s}\leq\bar{c}~\tilde{\sigma}_{*}^{2}N$
with high probability, where $\tilde{\sigma}^{2}_{*}=\max\tilde{\sigma}^{2}_{s}$. This completes the proof. ∎∎
To complete the proof of Lemma 3.5, we use this choice of $L$ and $\nu(\boldsymbol{Z})$ in the matrix Bernstein inequality to bound the probability that $\|\boldsymbol{Z}\|$ exceeds
$t:=c\max\left\{\sqrt{\tilde{\sigma}^{2}N\log N},(\log N)^{3/2}\right\}.$ (74)
To do so, We’ll examine various cases.
First, suppose that $(\log N)^{3/2}\leq\sqrt{\tilde{\sigma}^{2}N\log N}$ so that
$t=\sqrt{\tilde{\sigma}^{2}N\log N}.$
Substituting the values of $L$ and $\nu(\boldsymbol{Z})$ from Lemma A.1 and Lemma A.2 gives
$\nu(\boldsymbol{Z})+\frac{Lt}{3}$ $=\bar{c}\tilde{\sigma}^{2}N+\left(\sqrt{\max\{\tilde{\sigma}^{2},1\}\log N}\right)\left(\sqrt{\tilde{\sigma}^{2}N\log N}\right)$
$\leq\bar{c}\tilde{\sigma}^{2}N+\mathcal{O}\left(\tilde{\sigma}^{2}\log N\cdot\sqrt{N}\right)$
$=\mathcal{O}\left(\tilde{\sigma}^{2}N\right).$
It follows that there is some constant $\hat{c}>0$ such that
$\frac{t^{2}/2}{\nu(\boldsymbol{Z})+\frac{Lt}{3}}\geq\frac{c^{2}\tilde{\sigma}^{2}N\log N}{\hat{c}\tilde{\sigma}^{2}N}=\frac{c^{2}}{\hat{c}}\log N.$
Applying (68), we have
$\operatorname{Pr}(\|\boldsymbol{Z}\|\geq t)\leq(n+N)\left(\exp\left(-\frac{c^{2}}{\hat{c}}\log N\right)\right)\leq 2N^{-5}$
for $c$ sufficiently large.
Next, suppose that
$t=(\log N)^{3/2}\geq\max\left\{\sqrt{\tilde{\sigma}^{2}N\log N},(\log N)^{3/2}\right\}.$
Then
$\nu(\boldsymbol{Z})+\frac{Lt}{3}=\mathcal{O}\left((\log N)^{2}\right)$
and
$\frac{t^{2}/2}{\nu(\boldsymbol{Z})+\frac{Lt}{3}}\geq\frac{c^{2}(\log N)^{3}}{\hat{c}(\log N)^{2}}=\hat{c}\log N.$
Again, it follows that
$\operatorname{Pr}(\|\boldsymbol{Z}\|\geq t)\leq 2N^{-5}$
for sufficiently large $c$ in the choice of $t$. This completes the proof of Lemma 3.5. ∎